### v3

Abstract
Calibration allows predictions to be reliably interpreted as probabilities by decision makers. We propose a decision-theoretic calibration error, the Calibration Decision Loss (CDL), defined as the maximum improvement in decision payoff obtained by calibrating the predictions, where the maximum is over all payoff-bounded decision tasks. Vanishing CDL guarantees the payoff loss from miscalibration vanishes simultaneously for all downstream decision tasks. We show separations between CDL and existing calibration error metrics, including the most well-studied metric Expected Calibration Error (ECE). Our main technical contribution is a new efficient algorithm for online calibration that achieves near-optimal $O(\frac{\log T}{\sqrt{T}})$ expected CDL, bypassing the $\Omega(T^{-0.472})$ lower bound for ECE by Qiao and Valiant (2021).
1 Introduction
Given a sequence of predictions indicating, say, the chances of rain each day for a period of $T$ days, an intuitive way to assess the quality of these predictions is to check for calibration: for example, among the days predicted to have a 60% chance of rain, is the fraction of rainy days indeed 60%? Formally, suppose a prediction of $q_{i}\in[0,1]$ is received on $n_{i}$ days, and among the $n_{i}$ days, $m_{i}$ days are actually rainy. Calibration, a notion that originated from the forecasting literature (Dawid, 1982), requires the relationship $m_{i}=q_{i}n_{i}$ to hold for every prediction value $q_{i}$. The most well-studied calibration error metric is the Expected Calibration Error (ECE), defined as $\frac{1}{T}\sum_{i}|m_{i}-q_{i}n_{i}|$, or equivalently written as the average prediction bias $\frac{1}{T}\sum_{i}n_{i}|q_{i}-\frac{m_{i}}{n_{i}}|$, where $|q_{i}-\frac{m_{i}}{n_{i}}|$ is the absolute bias when $q_{i}$ is predicted. This gives a well-defined calibration error metric: a predictor is perfectly calibrated if and only if its ECE is zero.
In this work, we focus on the value of calibration for decision making. In the economics literature, the value of predictions is quantified by the payoff from decision making (Blackwell, 1951). Consider a decision maker who needs to choose an action $a\in A$ to maximize their decision payoff $U(a,\theta)$ , a (arbitrary) function of the action $a$ (e.g. take an umbrella or not) and the true binary state $\theta\in\{0,1\}$ (e.g. rainy or not). When the true state $\theta$ is unknown and only a prediction $p\in[0,1]$ is given, the decision maker can trust the prediction and take the action $a$ that maximizes the expected payoff $\operatorname{\mathbb{E}}_{\theta\sim p}[U(a,\theta)]$ , where “trust” means the decision maker assumes that the distribution of the state is indeed as predicted ($\Pr[\theta=1]=p$). Thus, the prediction manifests its value in the payoff received by the decision maker who best responds to the prediction as if the predictions were correct.
In the following examples, we explain how calibrated predictions are trustworthy for all downstream decision-makers, regardless of their specific decision tasks. Example 1.1 shows a miscalibrated predictor and Example 1.2 shows a calibrated predictor.
Example 1.1.
Among the $T$ samples of (prediction, state), $\frac{T}{2}$ predictions are $0.4$ and the other half are $0.6$. When the prediction is $0.4$, the actual empirical frequency of state $1$ is $0.2$; when the prediction is $0.6$, the actual empirical frequency of state $1$ is $0.8$.
Prediction $p_{t}$ 0.4 0.6 0.4 0.4 0.6 0.4 0.6 0.6 0.4 0.6
State $\theta_{t}$ 0 1 1 0 1 0 1 0 0 1
After observing enough samples of (prediction, state), the decision maker will figure out the “meaning” of each prediction: $0.4$ actually means a $0.2$ probability of state being $1$, and $0.6$ actually means $0.8$. Based on this observation, the decision maker will no longer trust the original prediction values when making decisions. Instead, if we calibrate the predictions by changing them to the actual empirical frequencies (e.g. $0.4\to 0.2$, and $0.6\to 0.8$, see Example 1.2), the decision maker can now trust the predictions by interpreting them as the true probabilities when making decisions.
Example 1.2.
Among the $T$ samples of (prediction, state), $\frac{T}{2}$ predictions are $0.2$ and the other half are $0.8$. When the prediction is $0.2$, the actual empirical frequency of state $1$ is indeed $0.2$; when the prediction is $0.8$, the actual empirical frequency of state $1$ is indeed $0.8$.
Prediction $p_{t}$ 0.2 0.8 0.2 0.2 0.8 0.2 0.8 0.8 0.2 0.8
State $\theta_{t}$ 0 1 1 0 1 0 1 0 0 1
1.1 The Decision Loss from Miscalibration
The examples above seem to suggest that calibration is a valuable property of a predictor for downstream decision makers. Is there a way to precisely quantify this value? Does the widely used calibration error ECE give the right quantification? Our main result reveals a fundamental gap between ECE and the value of calibration for decision making. In fact, ECE overestimates the decision loss from miscalibration.
To provide some intuition for this gap, let us consider the miscalibrated predictor in Example 1.1 and a decision task defined as follows, where we show that ECE does not give the payoff loss due to miscalibration.
The decision task has binary states $\Theta=\{0,1\}$ (not rainy or rainy) and actions $A=\{0,1\}$, where $1$ stands for taking an umbrella and 00 for not. The payoff is $1$ if the decision maker’s action matches the state and 00 otherwise. Best responding to the prediction, the decision maker chooses to bring an umbrella when the predicted chance of rain is at least half ($p_{t}\geq 0.5$). The example is visually explained in Figure 1.
[…]
Figure 1: In this example, ECE overestimates the decision loss from miscalibration. The plot visualizes the predictions in $[0,1]$. The best-response decision rule changes action at threshold $\nicefrac{{1}}{{2}}$ (red). The ground truth probability takes values from $0.2$ and $0.8$. The miscalibrated predictor predicts $0.4$ (blue) when the ground truth is $0.2$, and $0.6$ (blue) when the ground truth is $0.8$. Miscalibration induces no loss to the decision maker, since both the ground truth probability and the prediction recommends the same action.
The average prediction bias, ECE, of the (miscalibrated) predictor in Example 1.1 is $0.2$. However, the decision maker’s payoff loss from this miscalibration is actually 00. This is because the decision maker would take the same actions even if the predictions are recalibrated as in Example 1.2. Indeed, a prediction $0.4$ and the recalibrated version $0.2$ are both below the decision threshold of $0.5$, leading to action 00. Similarly, a prediction $0.6$ and the recalibrated version $0.8$ both lead to action $1$.
In this example, for the particular decision maker, the loss caused by miscalibration is zero, i.e. recalibrating the predictions provides zero value. In contrast, ECE is $0.2>0$, overestimating the decision loss from miscalibration. While this example only considers a specific decision task with two actions, we ask the natural question for all decision tasks with arbitrarily many actions - does ECE simultaneously and significantly overestimate the decision loss for every downstream task?
To answer this question, we propose a calibration error for decision making, the Calibration Decision Loss (CDL), which quantifies the worst-case payoff loss caused by miscalibration of a predictor. Our main result demonstrates a fundamental separation between CDL and existing error metrics including ECE. Inspired by this separation, we give a new efficient algorithm for online prediction that minimizes CDL at a faster rate than what is possible for ECE (see Section 1.2 for more details).
Our definition of CDL can be decomposed into two steps. First, for a fixed decision task, miscalibration incurs a loss to the decision maker, which we call Calibration Fixed Decision Loss (CFDL). Consider a decision maker who best responds to a miscalibrated predictor as in Example 1.1. We define CFDL as the payoff increase (averaged over $T$ rounds) that the decision maker could have achieved if we had calibrated the predictions as in Example 1.2. Second, since perfect calibration guarantees trustworthiness for all decision makers, we measure the calibration error by defining CDL as the worst-case (i.e., maximum) CFDL over all decision tasks with payoffs bounded in $[0,1]$. Here we restrict the payoffs to be bounded solely for the purpose of normalization: the CFDL of a decision task scales proportionally if we multiply the payoffs by any positive constant, and it remains the same if any constant is added to the payoffs. Beyond that, we make no additional restrictions on the decision tasks. In particular, we allow each decision task to have an arbitrarily large action space $A$, and the CDL is the maximum CFDL over all such tasks. Thus, vanishing CDL implies that the payoff loss from miscalibration vanishes for all decision tasks.
Similar to ECE, our CDL is a well-defined calibration error metric, with the basic property that a sequence of predictions is perfectly calibrated if and only if its CDL is zero, as implied by Foster and Vohra (1997). [^2] In the binary decision task in Figure 1, the CFDL of the miscalibrated predictor (Example 1.1) is 00. However, other decision tasks exist where the miscalibrated predictor has positive CFDL, leading to a non-zero CDL. However, we show that CDL is different from, and is not even a constant factor approximation of existing calibration error metrics, including ECE, the $K_{2}$ calibration error (i.e. the expected squared prediction bias), and the smooth calibration error (Błasiok et al., 2023a). In particular, we observe that ECE is often a significant overestimation of CDL. As a strong demonstration of this observation, we give an efficient algorithm for online prediction that minimizes CDL at a near-optimal rate of $O(T^{-1/2}\log T)$, which is faster than what is possible for ECE: it surpasses the $\Omega(T^{-0.472})$ lower bound for ECE by Qiao and Valiant (2021). We discuss this result with more details in Section 1.2 below.
1.2 Online Calibration
Let us now consider the algorithmic task of making sequential predictions for $T$ rounds. The goal is to achieve asymptotic calibration, meaning that the predictions are arbitrarily “close” to perfect calibration for sufficiently large $T$, with “closeness” measured by some calibration error. Though it may seem impossible, a classic and remarkable result by Foster and Vohra (1998) shows that asymptotic calibration can be achieved without any knowledge of what the state will be in each round.
Concretely, Foster and Vohra (1998) consider the following online binary prediction setting. The algorithm (i.e., predictor) interacts with an adversary for $T$ rounds. In each round $t=1,\ldots,T$, the predictor makes a prediction $p_{t}\in[0,1]$ and the adversary reveals the true state $\theta_{t}\in\{0,1\}$. The only knowledge that the predictor can use to produce $p_{t}$ is the history $(p_{1},\theta_{1},\ldots,p_{t-1},\theta_{t-1})$, which may contain zero information about the new state $\theta_{t}$. The adversary, on the other hand, can choose $\theta_{t}$ arbitrarily, based on the history and the predictor’s strategy, so as to maximize the calibration error of the predictions.
Foster and Vohra (1998) focus on ECE as the metric for calibration error. They prove the existence of a randomized algorithm that guarantees $O(T^{-1/3})$ expected ECE, where the expectation is over the randomness of the algorithm. Their use of a randomized algorithm breaks the impossibility of asymptotic calibration if a deterministic prediction strategy were used. To see this impossibility, if the algorithm predicts deterministically, the adversary can pick $\theta_{t}=0$ whenever the prediction $p_{t}$ is above $0.5$, and $\theta_{t}=1$ otherwise, yielding $\textsc{ECE}\geq 1/2$. In contrast, the predictor in Foster and Vohra (1998) draws $p_{t}$ randomly from a distribution. The adversary, knowing the predictor’s strategy, can choose $\theta_{t}$ based on this distribution, but it can no longer choose $\theta_{t}$ based on the realized value of $p_{t}$ drawn from the distribution.
This remarkable first result of Foster and Vohra (1998) opened up the literature of online asymptotic calibration, with the focus of error metric mainly on ECE. On the upperbound side, subsequent work constructs polynomial-time algorithms for ECE minimization (e.g. Foster, 1999; Foster and Hart, 2021). On the lowerbound side, a recent breakthrough of Qiao and Valiant (2021) proves an $\Omega(T^{-0.472})$ lower bound for ECE, improving over the long-known natural $\Omega(T^{-1/2})$ bound. [^3] Subsequent to our work, Dagan et al. (2024) shows there exists a sequential prediction strategy that achieves strictly lower than $T^{-\nicefrac{{1}}{{3}}}$ expected ECE. They also give an improved lowerbound of ECE over $\Omega(T^{-0.472})$ in Qiao and Valiant (2021).
It is unclear, however, that for decision making, ECE is the calibration error metric we wish to optimize. The economic value of making good predictions lies in helping downstream decision makers achieve better payoff. If the goal of calibration is to generate trustworthy predictions that induce no payoff loss to all decision makers, what we should really minimize is CDL, rather than ECE.
Prior to our work, nothing was known about CDL minimization beyond what is implied by minimizing ECE. Specifically, the relationship $\textsc{CDL}\leq 2\,\textsc{ECE}$ can be inferred from Kleinberg et al. (2023). Therefore, any ECE minimization algorithm also guarantees that CDL vanishes at the same rate (i.e., up to a factor of $2$). The interesting question is to go beyond this implication: can we achieve a better error rate for CDL than that is possible for ECE, surpassing the $\Omega(T^{-0.472})$ lower bound?
Our work gives a positive answer. We give an efficient randomized algorithm that guarantees $O(T^{-1/2}\log T)$ expected CDL (Theorem 5.1). This error rate surpasses the $\Omega(T^{-0.472})$ lowerbound for ECE and is optimal up to a logarithmic factor: there is a natural $\Omega(T^{-1/2})$ lower bound for CDL minimization from drawing each state $\theta_{t}$ independently and uniformly from $\{0,1\}$. [^4] Consider the decision task with two actions ($A=\{0,1\}$) and the 0-1 payoff function $U(a,\theta)=\mathbb{I}\left[a=\theta\right]$. When the states are independent and unbiased coin flips, the total payoff drops below $T/2-\Omega(\sqrt{T})$ with constant probability, inducing $\Omega(T^{-1/2})$ CFDL. By definition, our algorithm guarantees that for every downstream decision task, the payoff loss from miscalibration (CFDL) vanishes at the same near-optimal rate simultaneously.
Our work establishes a comprehensive understanding of the optimal rate for CDL, compared to the significant gaps in the current best upper and lower bounds for the other two recently popular metrics in online asymptotic calibration (see Section 1.5 for more details):
  • ECE: $O(T^{-1/3})$ and $\Omega(T^{-0.472})$ (Foster and Vohra, 1998; Qiao and Valiant, 2021);
  • Distance to calibration (Błasiok et al., 2023a): $O(T^{-1/2})$ and $\Omega(T^{-2/3})$ (Qiao and Zheng, 2024; Arunachaleswaran et al., 2024).
Certainly, our result would be impossible without a separation between CDL and ECE. This separation has been hinted in our examples earlier, and we will give a more in-depth explanation in Section 1.3.2 with an overview of the techniques in this paper.
Our result implies a substantial strengthening of an independent result by Roth and Shi (2024) on swap regret minimization. Swap regret minimization has been studied extensively in the literature of online learning (Hart and Mas-Colell, 2000; Blum and Mansour, 2007). Consider a fixed decision task repeated for $T$ rounds. The swap regret is the payoff increase (averaged over $T$ rounds) when the decision maker is allowed to apply a mapping $\sigma:A\to A$ and swap each action to another action in hindsight. The swap regret is stronger than the more commonly studied external regret, which only allows the decision maker to swap each action to a fixed action in hindsight. Foster and Vohra (1999) have observed that a predictor is perfectly calibrated if and only if it guarantees no swap regret for every downstream decision maker who trusts the prediction. Roth and Shi (2024) propose an algorithm that achieves $O(\frac{|A|}{\sqrt{T}})$ swap regret for every downstream decision maker, where $|A|$ is the number of actions.
In our paper, our algorithm guarantees the same near-optimal $O(\frac{\log T}{\sqrt{T}})$ swap regret for every downstream decision maker simultaneously, removing the dependence on the number $|A|$ of actions in Roth and Shi (2024). This directly follows from the fact that swap regret is upperbounded by CFDL for a fixed a decision task. To see this, first recall that CFDL is the improvement in payoff when predictions are empirically calibrated to their conditional empirical frequencies. While empirically calibrated predictions may suggest different actions to the decision maker, this improvement can be equivalently written as the regret when, in hindsight, the decision maker is allowed to swap actions whenever they receive a different prediction. Thus, CFDL is stronger since the modification rule is finer-grained than that of the swap regret. If two predictions suggest the same best-response action to a decision maker, the modification rule in CFDL allows the decision maker to swap the actions differently, while the swap regret does not.
The U-calibration error UCal, introduced by Kleinberg et al. (2023), is closely related to our CDL, despite being qualitatively weaker (i.e. UCal being zero is necessary but insufficient for calibration). Both errors are defined as the maximum payoff increase over decision tasks: CDL is the maximum payoff increase from calibrating the predictions; whereas U-calibration is the maximum payoff increase (a.k.a. external regret) from changing every prediction $p_{t}$ to the best fixed prediction (i.e., predicting the overall average $(\theta_{1}+\cdots+\theta_{T})/T$ every round). [^5] For a fixed decision task, the external regret may be negative, but the U-calibration error is non-negative because it is the maximum over all payoff-bounded decision tasks (including the degenerate task where the payoff is always zero). Therefore, the U-calibration error lowerbounds CDL. To see why UCal is not a well-defined calibration error, consider the miscalibrated predictor such that $(p_{t},\theta_{t})=(1/4,0)$ for $T/2$ rounds, and $(p_{t},\theta_{t})=(3/4,1)$ for the remaining $T/2$ rounds. The miscalibrated predictor always induce better or equal payoff compared to the best fixed prediction $1/2$, giving zero U-calibration error. See a detailed discussion in Example 4.5.
While our result appears very similar to Kleinberg et al. (2023), we use substantially different techniques since our CDL is qualitatively stronger. Kleinberg et al. (2023) give a randomized algorithm that guarantees an optimal $O(T^{-1/2})$ expected U-calibration error. Their prediction algorithm ensures that each best-responding decision maker takes actions as if they are locally running the classic Hedge algorithm (see Arora et al., 2012), which is specific to the external regret. Thus, their algorithm does not extend to our CFDL.
1.3 Results Overview
We discuss the connections between CDL and other calibration error metrics in Section 1.3.1. We describe our technical ideas behind our main result for online calibration in Section 1.3.2.
1.3.1 Properties of CDL
We propose CDL as a calibration error metric and study its properties. In Section 3.3, we show CDL can be computed in time polynomial in the size of the prediction space $|Q|$ and the state space $|\Theta|$ by solving a linear program, which holds for general non-binary state space.
In Section 4, we discuss the separation between CDL and other calibration error metrics, including ECE, the $K_{2}$ calibration error, and the smooth calibration error. We consider the average calibration error calculated on $T$ samples of the prediction and the state. Our results show previous results on minimizing $K_{2}$ and smCal do not directly apply. We also show the separation between CDL and the U-calibration error UCal in Kleinberg et al. (2023), where $\textsc{UCal}=0$ is necessary but not sufficient for the predictions to be calibrated.
  • ECE is polynomially related to CDL.
    $\textsc{ECE}^{2}\leq$ $\textsc{CDL}\leq 2\textsc{ECE}$
    We give examples where inequalities are asymptotically tight. In fact, the lower bound is attained by the same example in Section 1.3.2.
  • $K_{2}$, defined as average squared bias, is polynomially related to CDL.
    $K_{2}\leq$ $\textsc{CDL}\leq 2\sqrt{K_{2}}.$
    We give examples where inequalities are asymptotically tight. There exists an online algorithm that achieves $\tilde{O}(\frac{1}{\sqrt{T}})$ $K_{2}$ calibration error (Roth, 2022).
  • The smooth calibration error is not polynomially related to CDL, where we give examples.
  • The U-calibration error lowerbounds CDL, but is not polynomially related.
1.3.2 Online CDL Minimization: Technical Overview
The key idea behind our $O(\frac{\log T}{\sqrt{T}})$ guarantee for CDL comes from the observation that the CDL can often be significantly smaller than ECE, despite their linear relationship in the worst-case. This allows us to bypass the $\Omega(\frac{T^{0.528}}{T})$ lower bound for ECE. Here we describe a typical example where the CDL is much smaller than ECE. Based on the intuition from this example, we establish a general lemma (Lemma 1.3) which plays a crucial role in our analysis.
Let us discretize the prediction space by restricting our predictor to only produce predictions in a finite set $Q:=\{q_{1},\ldots,q_{m}\}\subseteq[0,1]$, where $q_{i}=i/m$ for $i=1,\ldots,m$. We view each $q_{i}$ as a “bucket”, and thus our predictor makes predictions that fall into these buckets. For a sequence of $T$ predictions $\bm{p}=(p_{1},\ldots,p_{T})\in Q^{T}$ made by our predictor and the corresponding true states $\bm{\theta}=(\theta_{1},\ldots,\theta_{T})\in\{0,1\}^{T}$, let $n_{i}$ denote the number of predictions in bucket $i$:
$n_{i}:=\sum_{t=1}^{T}\mathbb{I}\left[p_{t}=q_{i}\right],$ (1)
and let $\widehat{q}_{i}$ denote the empirical average of the true outcomes corresponding to the $n_{i}$ predictions:
$\widehat{q}_{i}:=\frac{1}{n_{i}}\sum_{t=1}^{T}\theta_{t}\mathbb{I}\left[p_{t}=q_{i}\right].$ (2)
ECE can be computed as
$\textsc{ECE}(\bm{p},\bm{\theta})=\sum_{i=1}^{m}n_{i}|q_{i}-\widehat{q}_{i}|.$ (3)
Assuming $\sqrt{T}$ is an integer, let us choose $m=\sqrt{T}$. Assume for simplicity that each bucket contains the same number of predictions: $n_{i}=\sqrt{T}$ for every $i=1,\ldots,m$. A typical guarantee one can often obtain using existing online learning techniques is the following bound on the deviation between $\widehat{q}_{i}$ and $q_{i}$:
$|\widehat{q}_{i}-q_{i}|\lesssim 1/\sqrt{n_{i}}=T^{-1/4}.$
Thus, let us assume $|\widehat{q}_{i}-q_{i}|=T^{-1/4}$ for $i=1,\ldots,m$ for simplicity. Plugging this into (3), we get $\textsc{ECE}(\bm{p},\bm{\theta})=T^{-1/4}$. This is even worse than the $\textsc{ECE}=O(T^{-1/3})$ guarantee by the algorithm of Foster and Vohra (1998), mainly because here we discretize the prediction space into $m=\sqrt{T}$ buckets, but the optimal choice would be $m\approx T^{1/3}$.
However, the CDL in this example is much smaller: in fact we have $\textsc{CDL}\leq O(T^{-1/2})$. To prove this fact, let us first consider a specific decision task with action space $A=[0,1]$, where the payoff is given by
$U(a,\theta)=\mathbb{I}\left[\theta=\mathbb{I}\left[a>0.5\right]\right]=\begin{cases}\mathbb{I}\left[\theta=0\right],&\text{if $a\leq 0.5$};\\ \mathbb{I}\left[\theta=1\right],&\text{if $a>0.5$}.\end{cases}$
Here, actions in the interval $[0,0.5]$ are equivalent in terms of their payoffs, and similarly for actions in $(0.5,1]$. Therefore, the best response strategy is not unique. The best response strategy which we focus on here is the identity function: $a^{*}(p)=p$. The calibration fixed decision loss (CFDL) on decision task $U$ is defined as the improvement in payoff after the decision maker calibrates the predictor:
$\textsc{CFDL}_{U}(\bm{p},\bm{\theta})=\frac{1}{T}\sum_{t}\bigg{[}U(\sigma(p_{t}),\theta_{t})-U(p_{t},\theta_{t})\bigg{]},$ (4)
where $\sigma$ is the mapping that swaps each prediction to the empirical frequency on samples.
We can decompose the calibration loss bucket-wise as follows:
$\textsc{CFDL}_{U}(\bm{a},\bm{\theta})$
[…]
If $q_{i}$ and $\widehat{q}_{i}$ belong to the same half of the interval $[0,1]$, i.e., $q_{i},\widehat{q}_{i}\in[0,0.5]$ or $q_{i},\widehat{q}_{i}\in(0.5,1]$, we clearly have
$\operatorname{\mathbb{E}}_{\theta\sim\widehat{q}_{i}}[U(\widehat{q}_{i},\theta)-U(q_{i},\theta)]=0.$ (6)
If $q_{i}$ and $\widehat{q}_{i}$ belong to different halves of the interval, a simple calculation gives
$\operatorname{\mathbb{E}}_{\theta\sim\widehat{q}_{i}}[U(\widehat{q}_{i},\theta)-U(q_{i},\theta)]=\max(\widehat{q}_{i},1-\widehat{q}_{i})-\min(\widehat{q}_{i},1-\widehat{q}_{i})=2|\widehat{q}_{i}-0.5|\leq 2|q_{i}-\widehat{q}_{i}|.$ (7)
Moreover, if $q_{i}$ and $\widehat{q}_{i}$ belong to different halves of the interval, by our assumption of $|\widehat{q}_{i}-q_{i}|=T^{-1/4}$, we have $|q_{i}-0.5|\leq T^{-1/4}$. Plugging (6) and (7) into (5), we get
$\textsc{CFDL}_{U}(\bm{a},\bm{\theta})\leq 2\sum_{i=1}^{m}n_{i}|q_{i}-\widehat{q}_{i}|\mathbb{I}\left[|q_{i}-0.5|\leq T^{-1/4}\right].$ (8)
Note that the number of $i$’s satisfying $|q_{i}-0.5|\leq T^{-1/4}$ is $O(T^{1/4})$. Thus, by our assumption of $n_{i}=\sqrt{T},|q_{i}-\widehat{q}_{i}|=T^{-1/4}$, we get $\textsc{CFDL}_{U}(\bm{a},\bm{\theta})=O(T^{-1/2})$.
While this bound on CFDL is for a specific decision task, we can extend it to every payoff-bounded decision task and get $\textsc{CDL}(\bm{p},\bm{\theta})=O(T^{-1/2})$. This follows from a result of Li ⓡ al. (2022) showing that any payoff-bounded decision task can be expressed as a convex combination of problems with V-shaped payoffs (see Section 3.2 for more details). Such decision tasks are all very similar to the one we consider here, and we can similarly obtain an $O(T^{-1/2})$ bound for the calibration loss for each of them. This reduction to V-shaped payoffs also played a crucial role in the $O(T^{-1/2})$ U-calibration guarantee of Kleinberg et al. (2023).
In our example above, we make the simplifying assumption that the number of predictions in each bucket is the same. However, the $O(T^{-1/2})$ bound on the CDL holds without this assumption while only losing a poly-logarithmic factor. That is, we have the following lemma:
Lemma 1.3 (Informal special case of Lemma 5.2 with $\alpha=O(1),\beta=0,\mathcal{D}(\bm{p},\bm{\theta})=0$).
Let $T,m$ be positive integers satisfying $m=\Theta(\sqrt{T})$. Define $Q=\{q_{1},\ldots,q_{m}\}\subseteq[0,1]$ where $q_{i}=i/m$ for every $i=1,\ldots,m$. Given a sequence of predictions $\bm{p}=(p_{1},\ldots,p_{T})\in Q^{T}$ and realized states $\bm{\theta}=(\theta_{1},\ldots,\theta_{T})\in\{0,1\}^{T}$, define $n_{i}$ and $\widehat{q}_{i}$ as in (1) and (2).
Assume $|\widehat{q}_{i}-q_{i}|\leq O(\frac{1}{\sqrt{n_{i}}})$ for every $i=1,\ldots,m$. Then
$\textsc{CDL}(\bm{p},\bm{\theta})\leq\widetilde{O}(T^{-1/2}).$
The intuition behind the lemma can be understood by analyzing the contribution of each bucket to CDL. In ECE, as expressed in Equation 3, the bias $n_{i}|q_{i}-\widehat{q}_{i}|$ in every bucket contributes to the sum, but in our example above, only a minority of the buckets make positive contribution to the CDL, as shown in (8). In general, for any specific decision task with V-shaped payoff, we show that the contribution to CDL from many buckets $i$ is significantly less than the bias $n_{i}|q_{i}-\widehat{q}_{i}|$. This allows us to prove the upper bound on CDL in Lemma 1.3, which would not hold if CDL were replaced by ECE.
Given Lemma 1.3, it remains to show that the guarantee $|\widehat{q}_{i}-q_{i}|\leq O(1/\sqrt{n}_{i})$ can indeed be (approximately) achieved in the online binary prediction setting. We use the result from Noarov et al. (2023) which, as stated, shows an efficient algorithm that gives us a bound only on the maximum of the expectation $\max_{i}({\mathbf{E}}\left[|\widehat{q}_{i}-q_{i}|\right]-O(1/\sqrt{n}_{i}))$, where the expectation is over the randomness in the algorithm. We refine their analysis and give a bound on the expectation of the maximum ${\mathbf{E}}\left[\max_{i}(|\widehat{q}_{i}-q_{i}|-O(1/\sqrt{n}_{i}))\right]$ (Lemma 5.5). As we show in Lemma 5.2 (a generalized version of Lemma 1.3), this bound is sufficient for us to obtain $\operatorname{\mathbb{E}}[\textsc{CDL}(\bm{p},\bm{\theta})]\leq O(\frac{\log T}{\sqrt{T}})$. We also note that a simpler minimax proof, similar to the one by Hart (2022), also allows us to show the existence of a randomized algorithm that approximately guarantees $|\widehat{q}_{i}-q_{i}|\leq O(1/\sqrt{n}_{i})$ and thus, by our Lemma 5.2 again, achieves $\operatorname{\mathbb{E}}[\textsc{CDL}(\bm{p},\bm{\theta})]\leq O(\frac{\log T}{\sqrt{T}})$ (see Appendix A). However, this proof does not come with an explicit construction or any efficiency guarantee.
1.4 Paper Organization
We introduce the preliminaries in Section 2, including popular measures of the calibration error, decision making and swap regret, and the online binary calibration problem. We introduce Maximum Calibration Loss (CDL) in Section 3 and discuss its alternative formulation using Bregman divergences and its approximation via V-shaped Bregman divergences, which will be useful to establish our main result. Section 4 discusses the connection between CDL and other calibration error metrics. In Section 5, we present our main result, an efficient online binary prediction algorithm that guarantees $O(\frac{\log T}{\sqrt{T}})$ CDL. The key technical idea behind this result is a lemma (Lemma 5.2) we prove in Section 5.1 which allows us to attribute CDL to bucket-wise biases. Additionally, we give a non-constructive but simpler minimax proof of the $O(\frac{\log T}{\sqrt{T}})$ CDL guarantee in Appendix A. This simpler proof also crucially relies on our key technical lemma (Lemma 5.2) in Section 5.1.
1.5 Related Work
1.5.1 Calibration Error Metrics
While perfect calibration has an intuitive and clear definition, it is a non-trivial and subtle question to meaningfully quantify the calibration error of predictions that are not perfectly calibrated. ECE is one of the most popular calibration measures, but it lacks continuity: slightly perturbing perfectly calibrated predictions can significantly increase ECE. To address this issue, recently Błasiok et al. (2023a) developed a theory of consistent calibration measures by introducing the distance to calibration as a central notion. This theory has facilitated rigorous explanations of an interesting empirical phenomenon called “calibration out of the box” in deep learning (Błasiok et al., 2023b, 2024).
As a relaxation of calibration, Kleinberg et al. (2023) consider the utility of predictions to downstream decision makers and introduce U-calibration error, the maximum external regret over payoff-bounded decision tasks. Kleinberg et al. (2023) design an algorithm that achieves an optimal $O(\frac{1}{\sqrt{T}})$ U-calibration, which is a necessary but insufficient condition for calibration.
1.5.2 Online Calibration Algorithms
Recent research has made significant progress in proving upper and lower bounds on the optimal rate achievable for both ECE and the distance to calibration in online binary prediction, though significant gaps remain between the current best upper and lower bounds. For ECE minimization, Foster and Vohra (1998) design a randomized algorithm that achieves $O(T^{-2/3})$ expected ECE, which remains the best known upper bound. Existence proofs and constructions of such algorithms have been further explored in several subsequent works (Hart, 2022; Foster and Hart, 2021). A recent work (Qiao and Valiant, 2021) show a lower bound of $\Omega(\frac{T^{0.528}}{T})$ to online ECE minimization. For the distance to calibration, Qiao and Zheng (2024) give a non-constructive minimax-based proof for an $O(\frac{1}{\sqrt{T}})$ upper bound and an $\Omega(T^{-1/3})$ lower bound for the same problem. Soon afterwards, Arunachaleswaran et al. (2024) provide a construction of an efficient algorithm that achieves the $O(\frac{1}{\sqrt{T}})$ upper bound for the distance to calibration.
The literature on online regret minimization for all downstream decision makers is technically closest to our paper. Kleinberg et al. (2023) observe that the swap regret of actions for any payoff-bounded decision task is linearly upperbounded by ECE. This observation quantitatively justifies the qualitative equivalence between no swap regret and calibration in Foster and Vohra (1997). However, the $\Omega(\frac{T^{0.528}}{T})$ lowerbound presents a barrier in efficient swap regret minimization via ECE minimization. To overcome this barrier from ECE, several relaxations of CDL have been considered by recent works to achieve $\tilde{O}(\frac{1}{\sqrt{T}})$ regret rates in online binary prediction. Kleinberg et al. (2023) achieve the $O(\frac{1}{\sqrt{T}})$ external regret maximized over payoff-bounded tasks, a necessary but insufficient condition for asymptotic calibration. On the other hand, Roth and Shi (2024), an independent work of ours, show $\widetilde{O}(|A|\frac{1}{\sqrt{T}})$ swap regret bounds that depend additionally on the number of actions $|A|$ in the downstream decision task, which can be loose when a decision task has many or even infinitely many actions.
1.5.3 Omniprediction
Treating prediction and decision making as separate steps allows us to train a single predictor and use it to solve multiple decision tasks with different utility/loss functions. This separation of training and decision making is the idea behind omniprediction, introduced recently by Gopalan et al. (2022), where the goal is to train a single predictor that allows each downstream decision maker to incur comparable or smaller loss than any alternative decision rule from a benchmark class. Notions from the algorithmic fairness literature (e.g. multicalibration and multiaccuracy (Hebert-Johnson et al., 2018; Kim et al., 2019)) have been used to obtain omnipredictors in various online and offline (batch) settings (Gopalan et al., 2022, 2023a; Hu et al., 2023; Gopalan et al., 2023b, 2024; Garg et al., 2024; Noarov et al., 2023; Kim and Perdomo, 2023). Omniprediction allows better efficiency than training a different model from scratch for each decision task, and it also allows the predictor to be robust to changes in the loss function.
1.5.4 Swap Regret Minimization
Swap regret minimization algorithms have been studied extensively in the online learning literature (e.g. Hart and Mas-Colell, 2000, 2001; Blum and Mansour, 2007; Hart and Mas-Colell, 2013; Anagnostides et al., 2022). In game theory, the swap regret is known for its connection to correlated equilibrium. Foster and Vohra (1997) first show vanishing swap regret implies convergence to correlated equilibria. Recently, Peng and Rubinstein (2023); Dagan et al. (2023) prove a lowerbound on the swap regret, which is polynomial in the number of actions. Meanwhile, the calibration literature (Kleinberg et al., 2023; Noarov et al., 2023; Roth and Shi, 2024) differs from the swap regret minimization literature in two aspects: 1) it focuses on developing a robust strategy that minimizes swap regret simultaneously for all decision makers, and 2) it focuses on minimizing swap regret for the special payoff structure of decision tasks. As a special payoff structure, a decision task restricts the adversary to only be able to select a state. Equivalently, the adversary can select payoff from a low rank matrix with the same rank of the state space. Our result also implies the lowerbound on swap regret is strictly weaker when there exists a special low-rank constraint on payoff matrix. While the lowerbound on general swap regret minimization depends polynomially on the number of actions, swap regret minimization of decision tasks does not have such dependence given restricted state space.
1.5.5 Optimization of Scoring Rules
CDL is defined as the maximum swap regret over all decision tasks. Since the payoff in a decision task can be equivalently represented by proper scoring rules (see Section 2.3), the computation of CDL is an optimization problem of scoring rules. A recent literature (Li ⓡ al., 2022; Neyman et al., 2021; Hartline ⓡ al., 2023) studies the optimization of scoring rules, where Li ⓡ al. (2022) is the most relevant paper. Li ⓡ al. (2022) present two results that are helpful to our problem: 1) under their different optimization objective, the optimal scoring rule can be computed via linear programming, and 2) any bounded scoring rule can be decomposed into a linear combination of V-shaped scoring rules. Following their idea, we design a linear program that computes CDL in polynomial time. Our $O(\frac{\log T}{\sqrt{T}})$ CDL result also uses this linear decomposition of scoring rules (see Section 3.2).
2 Preliminaries
Throughout the paper, we use $p\in[0,1]$ to denote a prediction, and use $\theta$ to denote a binary state. A prediction $p$ can be viewed as a distribution over the state space $\Theta=\{0,1\}$, and we write $\theta\sim p$ when we sample $\theta$ from the Bernoulli distribution with mean $p$, i.e., $\Pr[\theta=1]=p$. For a real number $x$, we use $(x)_{+}$ or $[x]_{+}$ to denote $\max\{x,0\}$. We use $\mathbb{I}\left[\cdot\right]$ to denote the 0-1 indicator function: $\mathbb{I}\left[\text{statement}\right]=1$ if the statement is true, and $\mathbb{I}\left[\text{statement}\right]=0$ if the statement is false.
2.1 Measures of Calibration Error
In this section, we define empirical calibration on $T$ samples and calibration error metrics. Over $T$ samples, the predictions are restricted to fall in a finite space $Q=\{q_{i}\in[0,1]\}_{i}$. Let $n_{i}=\sum_{t}\mathbb{I}\left[p_{t}=q_{i}\right]$ be the count of prediction being $q_{i}$ in $T$ samples, and $\widehat{q}_{i}=\frac{\sum_{t}\mathbb{I}\left[p_{t}=q_{i}\right]\theta_{t}}{n_{i}}$ be the empirical distribution of the realized state conditioning on prediction is $q_{i}$.
Definition 2.1.
Given $T$ samples, predictor is empirically calibrated if for each prediction $q_{i}\in Q$, the prediction is consistent with its conditional empirical distribution, i.e. $q_{i}=\widehat{q}_{i}$.
Definition 2.2.
Given $T$ samples, we can empirically calibrate a prediction by swapping predictions to their conditional empirical frequencies, i.e. by applying swap mapping $\sigma^{*}$ to predictions with $\sigma^{*}(q_{i})=\widehat{q}_{i}$.
While there is only one natural and clear definition of perfect calibration, there are various metrics for measuring the calibration error. Foster and Vohra (1998) defines the expected calibration error (ECE) which is a measure of calibration error popularly used in the literature. ECE measures the average absolute distance between the prediction and the empirical distribution.
Definition 2.3 (ECE).
Given $T$ samples of predictions $\bm{p}=(p_{t})_{t\in[T]}$ and corresponding realizations $\bm{\theta}=(\theta_{t})_{t\in[T]}$ of states. ECE is
$\textsc{ECE}(\bm{p},\bm{\theta})=\frac{1}{T}\sum_{q_{i}\in Q}n_{i}|q_{i}-\widehat{q}_{i}|.$
An alternative metric is the $K_{2}$ calibration error, the average squared distance between a prediction and the empirical distribution. Theorem 2.23 shows it is possible to achieve $\tilde{O}(T^{-\nicefrac{{1}}{{2}}})$ worst-case expected $K_{2}$ error.
Definition 2.4 ($K_{2}$ calibration error).
Given $T$ samples of predictions $\bm{p}=(p_{t})_{t\in[T]}$ and corresponding realizations $\bm{\theta}=(\theta_{t})_{t\in[T]}$ of states, the $K_{2}$ calibration error is
$K_{2}(\bm{p},\bm{\theta})=\frac{1}{T}\sum_{q_{i}\in Q}n_{i}(q_{i}-\widehat{q}_{i})^{2}.$
In addition to ECE and $K_{2}$ calibration error, we also compare to the smooth calibration error introduced by Kakade and Foster (2008). Unlike ECE and $K_{2}$, the smooth calibration error is continuous in predictions.
Definition 2.5 (Smooth Calibration Error (Kakade and Foster, 2008)).
Given $T$ samples of predictions $\bm{p}=(p_{t})_{t\in[T]}$ and corresponding realizations $\bm{\theta}=(\theta_{t})_{t\in[T]}$ of states. The smooth calibration error is a supremum over the set $\Sigma$ of $1$-Lipschitz functions $\sigma:[0,1]\to[-1,1]$:
$\textsc{smCal}(\bm{p},\bm{\theta})=\frac{1}{T}\sup_{\sigma\in\Sigma}\sum_{t}\sigma(p_{t})(p_{t}-\theta_{t}).$
The definition above is equivalent to $K_{1}$ without the $1$-Lipschitz constraint on $\sigma$. Taking the following non-Lipschitz $\sigma$ yields ECE.
$\sigma(q_{i})=\begin{cases}1,&\text{if }q_{i}-\widehat{q}_{i}\geq 0;\\ -1,&\text{otherwise.}\end{cases}$
The smooth calibration error is polynomially related to the distance to calibration, which measures the absolute distance to the closest calibrated prediction.
Definition 2.6 (Distance to Calibration (Błasiok et al., 2023a)).
Given $T$ samples of predictions $\bm{p}=(p_{t})_{t\in[T]}$ and corresponding realizations $\bm{\theta}=(\theta_{t})_{t\in[T]}$ of states. The distance to calibration is
$\textsc{distCal}(\bm{p},\bm{\theta})=\frac{1}{T}\min_{\bm{\widehat{p}}:K_{1}(\bm{\widehat{p}},\bm{\theta})=0}\sum_{t}|p_{t}-\widehat{p}_{t}|.$
Lemma 2.7 (Błasiok et al. (2023a)).
Smooth calibration error smCal is polynomially related to distance to calibration distCal.
$\textsc{smCal}\leq\textsc{distCal}\leq\sqrt{32\,\textsc{smCal}}.$
2.2 Decision Making and Swap Regret
Before introducing predictions and calibration, we define decision making and swap regret. A decision task is defined by three components. Throughout the paper, we normalize the payoff in the decision task to $[0,1]$.
  • The agent takes action $a\in A$.
  • A payoff-relevant state $\theta\in\Theta$ realizes.
[…]
The decision is evaluated by the average performance when the decision maker repeatedly faces an identical decision task $(A,\Theta,U)$ in $T$ rounds. We define the swap regret the same as in Roth and Shi (2024).
Definition 2.8 (Swap Regret).
Given a sequence of $T$ actions $\bm{a}=(a_{t})_{t\in[T]}$ and realization $\bm{\theta}=(\theta_{t})_{t\in[T]}$ of states, and fix a decision task with payoff rule $U$, the swap regret of the decision maker is
$\textsc{Swap}_{U}(\bm{a},\bm{\theta})=\frac{1}{T}\max_{\sigma:A\to A}\sum_{t\in[T]}\bigg{[}U(\sigma(a_{t}),\theta_{t})-U(a_{t},\theta_{t})\bigg{]}.$
Kleinberg et al. (2023) aim to minimize the external regret for all decision tasks with bounded payoff. We define the external regret here for comparison.
Definition 2.9 (External Regret).
Given a sequence of $T$ actions $\bm{a}=(a_{t})_{t\in[T]}$ and realization $\bm{\theta}=(\theta_{t})_{t\in[T]}$ of states, and fix a decision task with payoff rule $U$, the external regret is calculated against the best fixed action,
$\textsc{Ext}_{U}(\bm{a},\bm{\theta})=\frac{1}{T}\max_{a\in A}\bigg{[}U(a,\theta_{t})-U(a_{t},\theta_{t})\bigg{]}.$ (9)
Kleinberg et al. (2023) define the U-calibration error, the maximum external regret when the decision maker best responds to the predictions. Note that no U-calibration error is necessary but insufficient for calibration.
Definition 2.10 (U-calibration Error).
Let $\bm{p}=(p_{t})_{t\in[T]}$ be a sequence of $T$ predictions and let $\bm{\theta}=(\theta_{t})_{t\in[T]}$ be the realization of states. Consider a decision maker with payoff function $U:A\times\{0,1\}\to[0,1]$ who best responds to the predictions by taking $a_{t}=a^{*}(p_{t})=\operatorname*{arg\,max}_{a\in A}{\mathbf{E}}_{\theta\sim p_{t}}\left[U(a,\theta)\right]$. Let $\bm{a}_{U}=(a_{1},\ldots,a_{T})$ be the vector of best-response actions. The U-calibration error is the maximum external regret over decision tasks with bounded payoff in $[0,1]$:
$\textsc{UCal}(\bm{p},\bm{\theta})=\sup_{U}\textsc{Ext}_{U}(\bm{a}_{U},\bm{\theta}),$
where the supremum is over all payoff functions $U:A\times\{0,1\}\to[0,1]$ with arbitrary action spaces $A$.
2.3 Payoff as Proper Scoring Rules
For ease of notation throughout the paper, we will write the payoff as a function of the prediction assuming the agent trusts the prediction and best responds to it. In this section, we introduce this payoff as a function of the prediction and show its equivalence to proper scoring rules.
When the agent is assisted with a sequence of predictions, the agent obtains payoff by acting in response to the prediction. If the prediction is calibrated, the best response $a^{*}$ is a function of the prediction:
$a^{*}(p)=\operatorname*{arg\,max}_{a\in A}{\mathbf{E}}_{\theta\sim p}\left[U(a,\theta)\right].$ (10)
The payoff is equivalently a function of the prediction. We present the payoff in a scoring rule $S:\Delta(\Theta)\times\Theta\to\mathbb{R}$, representing the utility directly as a function of the prediction.
Definition 2.11 (Scoring rule induced by decision task).
Given a decision task $(A,\Theta,U)$ and its corresponding best response function $a^{*}$, we define an induced scoring rule $S_{U}:\Delta(\Theta)\times\Theta\to\mathbb{R}$ such that for any prediction $p\in\Delta(\Theta)$ and state $\theta\in\Theta$,
$S_{U}(p,\theta)=U(a^{*}(p),\theta).$
That is, $S_{U}(p,\theta)$ is the payoff of a decision maker who chooses the best response $a^{*}(p)$ based on the prediction $p$.
Frongillo and Kash (2014) observe that scoring rules induced by a decision task are equivalent to the class of proper scoring rules (McCarthy, 1956). Proper scoring rules are defined such that the expected score is maximized when the prediction is the true distribution. Evaluated with a proper scoring rule, the predictor has an incentive to truthfully report the distribution in order to maximize expected score.
Definition 2.12 (Proper Scoring Rule).
A scoring rule $S$ is proper if for any $p^{\prime}\in\Delta(\Theta)$ that is not the true distribution $\theta$ is generated,
${\mathbf{E}}_{\theta\sim p}\left[S(p,\theta)\right]\geq{\mathbf{E}}_{\theta\sim p}\left[S(p^{\prime},\theta)\right].$
2.13 shows that 1) the scoring rule induced by a decision task is a proper scoring rule, and 2) any proper scoring rule measures the best-response payoff in a decision task.
Claim 2.13.
Proper scoring rules and payoffs in decision task are equivalent:
  1. 1.
    Any scoring rule $S_{U}$ induced by a decision task $(A,\Theta,U)$ is proper.
  2. 2.
    For any proper scoring rule $S$, there exists a decision task with payoff $U$, such that $U(a^{*}(p),\theta)=S(p,\theta)$ for every prediction $p$.
Proof.
2 is straight forward by defining $A=\Delta(\Theta)$, the action space as the prediction space.
To see 1, notice that the score is the best-response score with $S_{U}(p,\theta)=U(a^{*}(p),\theta)$. Hence, if the prediction is $p^{\prime}\neq p$ not the distribution where $\theta$ is drawn, the agent obtains a payoff weakly lower than the payoff by best responding to the true distribution.
${\mathbf{E}}_{\theta\sim p}\left[S(p^{\prime},\theta)\right]={\mathbf{E}}_{\theta\sim p}\left[U(a^{*}(p^{\prime}),\theta)\right]\leq{\mathbf{E}}_{\theta\sim p}\left[U(a^{*}(p),\theta)\right]={\mathbf{E}}_{\theta\sim p}\left[S(p,\theta)\right].\qed$
Following this equivalence between proper scoring rules and best-respond payoffs, we will introduce our results in a proper scoring rule $S$ instead of the decision payoff $U$.
2.4 Calibration Fixed Decision Loss
In the previous section, we assume that the agent trusts the prediction and best responds to it. If, however, the agent best responds to miscalibrated predictions, she suffers the Calibration Fixed Decision Loss (CFDL), the improvement in payoff when she empirically calibrates the predictions.
Definition 2.14 (CFDL).
Given $T$ samples of predictions $\bm{p}=(p_{t})_{t\in[T]}$ and corresponding realizations $\bm{\theta}=(\theta_{t})_{t\in[T]}$ of states, fixing a proper scoring rule $S$, the Calibration Fixed Decision Loss (CFDL) is
$\textsc{CFDL}_{S}=\frac{1}{T}\sum_{t}\bigg{[}S(\sigma^{*}(p_{t}),\theta_{t})-S(p_{t},\theta_{t})\bigg{]},$
where $\sigma^{*}$ is the swap mapping that empirically calibrates the predictor (see Definition 2.2).
The CFDL has an equivalent formalization as prediction swap regret, where the agent is allowed to swap predictions in hindsight. Note that the prediction swap regret is weakly stronger than the swap regret (Definition 2.8) on the same decision task.
Definition 2.15 (Prediction Swap Regret).
Given a sequence of $T$ predictions $\bm{p}=(p_{t})_{t\in[T]}$ and realization $\bm{\theta}=(\theta_{t})_{t\in[T]}$ of states, fixing a proper scoring rule $S$, the prediction swap regret is
$\textsc{PSwap}_{S}(\bm{p},\bm{\theta})=\max_{\sigma:\Delta(\Theta)\to A}\sum_{t\in[T]}\bigg{[}S(\sigma(p_{t}),\theta_{t})-S(p_{t},\theta_{t})\bigg{]},$
where $a^{*}$ is the best response to prediction as defined in Equation 10.
Proposition 2.16 shows CFDL is equivalent to prediction swap regret.
Proposition 2.16.
Given a sequence of $T$ predictions $\bm{p}=(p_{t})_{t\in[T]}$ and realization $\bm{\theta}=(\theta_{t})_{t\in[T]}$ of states, fixing a proper scoring rule $S$, CFDL equals the prediction swap regret:
$\textsc{CFDL}_{S}=\textsc{PSwap}_{S}.$
Proposition 2.16 follows directly from the properness of scoring rules. Conditioning on the prediction is $q_{i}$, the state follows the empirical distribution:
$\frac{1}{n_{i}}\sum_{t\in[T]}S(p,\theta_{t})\mathbb{I}\left[p_{t}=q_{i}\right]={\mathbf{E}}_{\theta\sim\widehat{q}_{i}}\left[S(p,\theta)\right].$
By properness of the scoring rule, predicting the empirical distribution maximizes the expected score (average score). Thus, the optimal swap function in Definition 2.15 is $\sigma^{*}$ which empirically calibrates the predictor.
The prediction swap regret is stronger than the swap regret in Definition 2.8. If an algorithm generates a sequence of predictions with low prediction swap regret, then the agent has low swap regret if they best respond to the predictions. To see this, notice that the modification rule in prediction swap regret has more power. The prediction swap regret allows the agent to modify the action conditioning on each prediction. If two predictions have the same best-response action, the swap regret does not allow the agent to apply different modification to the two predictions.
Claim 2.17.
Given a decision task with payoff rule $U$, the corresponding scoring rule is denoted $S$. If the agent best responds by taking $a_{t}=a^{*}(p_{t})$ at each round $t$,
$\textsc{Swap}_{U}(\bm{a},\bm{\theta})\leq\textsc{PSwap}_{S_{U}}$
Claim 2.18 (Kleinberg et al. (2023), Theorem 12).
 [^7] Kleinberg et al. (2023) assume that the output of the scoring rule $S$ is in $[-1,1]$, whereas we assume the output is in $[0,1]$. Thus, the constant $4$ in their bound translates to the constant $2$ here.
For any proper scoring rule $S:[0,1]\times\{0,1\}\to[0,1]$ and any sequences $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T},\bm{\theta}=(\theta_{1},\ldots,\theta_{T})\in\{0,1\}^{T}$, it holds that
$\textsc{CFDL}_{S}=\textsc{PSwap}_{S}(\bm{p},\bm{\theta})\leq 2\textsc{ECE}(\bm{p},\bm{\theta}).$
$K_{2}$ calibration error can be written as a special case of CFDL for quadratic scoring rule (a.k.a. squared loss). In fact, the prediction swap regret was first introduced in Foster and Vohra (1998) with the quadratic scoring rule $S(p,\theta)=1-(p-\theta)^{2}$.
Lemma 2.19.
Define the quadratic scoring rule $S_{2}(p,\theta)=1-(p-\theta)^{2}$. We have
$K_{2}=\textsc{CFDL}_{S_{2}}.$
ECE cannot be represented by the CFDL with any proper scoring rule
Lemma 2.20.
There does not exist a proper scoring rule $S$, such that for any sequence of predictions $\bm{p}$ and states $\bm{\theta}$,
$\textsc{ECE}(\bm{p},\bm{\theta})=\textsc{CFDL}_{S}(\bm{p},\bm{\theta}).$
Lemma 2.19 follows immediately from the definitions of $K_{2}$ and $S_{2}$, whereas Lemma 2.20 can be proved using the Bregman divergence characterization of proper scoring rules (see Proposition 3.6).
2.5 Online Binary Calibration
We focus on the classic online prediction problem studied by Foster and Vohra (1998). The goal is to generate calibrated predictions in the long run, even if the states are adversarially selected. Unless otherwise specified, we present results under the binary prediction setting, i.e. the state space is $\Theta=\{0,1\}$. In this setup, a predictor (algorithm) $F$ makes a prediction $p_{t}\in[0,1]$ at each round $t=1,2,\ldots$, and an adversary $A$ picks a binary state (outcome) $\theta_{t}\in\{0,1\}$. Both the prediction $p_{t}$ and the state $\theta_{t}$ can depend on the past history $h_{t-1}=(p_{1},\theta_{1},\dots,p_{t-1},\theta_{t-1})$, but they cannot depend on each other. That is, we can assume without loss of generality that the algorithm chooses $p_{t}$ and the adversary reveals $\theta_{t}$ simultaneously. The transcript $h_{t}=(p_{1},\theta_{1},\ldots,p_{t},\theta_{t})$ is a function of the strategies of the predictor $F$ and the adversary $A$. That is, $h_{t}=h_{t}(F,A)$. The predictions in $T$ rounds are evaluated by a calibration error metric.
We allow the predictor to be randomized, in which case we can view its strategy as a distribution ${\mathcal{F}}$ over deterministic strategies $F$. When we use an error metric, say MSR, to evaluate the predictions made by our predictor in $T$ rounds, our goal is to minimize the expected value of the error metric w.r.t. the worst-case adversary $A$, i.e., we want ${\mathbf{E}}_{F\sim{\mathcal{F}}}\left[\textsc{CDL}(h_{T}(F,A))\right]$ to be small for every adversary $A$.
Previous results for different calibration errors are listed here.
ECE
There exists a gap between the known upperbound and lowerbound for ECE.
Upperbound
Theorem 2.21 (Foster and Vohra (1998), see also Hart (2022); Foster and Hart (2021)).
There exists a randomized online binary prediction algorithm that guarantees $O(T^{-\nicefrac{{1}}{{3}}})$ expected ECE.
Lowerbound
Theorem 2.22 (Qiao and Valiant (2021)).
For any randomized online binary prediction algorithm, the ECE w.r.t. the worst-case adversary is $\Omega(T^{-0.472})$.
$K_{2}$ Calibration Error
As a variant of ECE, $K_{2}$ calibration error has an upperbound result from the literature.
Upperbound
Theorem 2.23 (Roth (2022)).
There exists a randomized online binary prediction algorithm that guarantees $O(T^{-\nicefrac{{1}}{{2}}}\log T)$ expected $K_{2}$ calibration error.
Smooth Calibration Error
Błasiok et al. (2023a) propose the smooth calibration error.
Upperbound
Qiao and Zheng (2024) prove the existence of an algorithm that achieves $O(\frac{1}{\sqrt{T}})$ distance to calibration, following which Arunachaleswaran et al. (2024) give a construction of such an algorithm. The result implies $O(\frac{1}{\sqrt{T}})$ smooth calibration error, since the distance to calibration upperbounds the smooth calibration error.
Theorem 2.24 (Qiao and Zheng (2024); Arunachaleswaran et al. (2024)).
There exists a randomized online binary prediction algorithm that guarantees $O(\frac{1}{\sqrt{T}})$ expected distance to calibration.
Lowerbound
Theorem 2.25 (Qiao and Zheng (2024)).
For any randomized online binary prediction algorithm, the smooth calibration error w.r.t. the worst-case adversary is $\Omega(T^{-\frac{2}{3}})$.
3 Calibration Decision Loss
In this section, we introduce our calibration error metric, the Calibration Decision Loss (CDL), meaning the decision loss resulted from miscalibration. In Sections 3.1 and 3.2, we discuss connections between CDL and Bregman divergences that will be useful for obtaining our main result in Section 5. We show an efficient algorithm for computing CDL in Section 3.3.
Definition 3.1 (CDL).
Given a sequence of $T$ predictions $\bm{p}=(p_{t})_{t\in[T]}$ and realization $\bm{\theta}=(\theta_{t})_{t\in[T]}$ of states, we define the Calibration Decision Loss (CDL) as
$\textsc{CDL}(\bm{p},\bm{\theta})=\sup_{S\in[0,1]}\textsc{CFDL}_{S}(\bm{p},\bm{\theta}),$
where the supremum is over all proper scoring rules $S$ with range bounded in $[0,1]$.
The CDL as defined above is equal to the maximum swap regret (Definition 2.8) of best responding decision makers with payoffs bounded in $[0,1]$:
Lemma 3.2.
Let $\bm{p}=(p_{t})_{t\in[T]}$ be a sequence of $T$ predictions and let $\bm{\theta}=(\theta_{t})_{t\in[T]}$ be the realization of states. Consider a decision maker with payoff function $U:A\times\{0,1\}\to[0,1]$ who best responds to the predictions by taking $a_{t}=a^{*}(p_{t})=\operatorname*{arg\,max}_{a\in A}{\mathbf{E}}_{\theta\sim p_{t}}\left[U(a,\theta)\right]$. Let $\bm{a}_{U}=(a_{1},\ldots,a_{T})$ be the vector of best-response actions. We have
$\textsc{CDL}(\bm{p},\bm{\theta})=\sup_{U}\textsc{Swap}_{U}(\bm{a}_{U},\bm{\theta}),$
where the supremum is over all payoff functions $U:A\times\{0,1\}\to[0,1]$ with arbitrary action spaces $A$.
Proof.
For any decision task with payoff function $U$, the prediction swap regret for the corresponding scoring rule $S$ is higher than the swap regret: $\textsc{Swap}_{U}(\bm{a}_{U},\bm{\theta})\leq\textsc{PSwap}_{S_{U}}(\bm{p},\bm{\theta})$ (2.17). Thus, $\textsc{CDL}\geq\sup_{U}\textsc{Swap}_{U}(\bm{a}_{U},\bm{\theta})$. On the other hand, for any scoring rule $S$, construct a decision task with payoff $U$ as in the proof of 2.13, by setting $A=[0,1]$. The resulting decision task has $\textsc{PSwap}_{S}=\textsc{Swap}_{U}$. Thus we have $\textsc{CDL}=\sup_{U}\textsc{Swap}_{U}(\bm{a}_{U},\bm{\theta})$. ∎
By definition, if CDL vanishes, then CFDL (and thus swap regret) also vanishes for every downstream agent.
Proposition 3.3.
If $\textsc{CDL}=\tilde{O}(\frac{1}{\sqrt{T}})$, then both $\textsc{CFDL}=\tilde{O}(\frac{1}{\sqrt{T}})$ and $\textsc{Swap}=\tilde{O}(\frac{1}{\sqrt{T}})$ simultaneously for every downstream decision task with payoff bounded in $[0,1]$.
3.1 Bregman Divergence Formulation
As an important preparation for our main result in Section 5, we show that the CFDL can be written in an equivalent form using the Bregman divergence (Proposition 3.6). This follows from the convex function formulation of proper scoring rules from McCarthy (1956); Savage (1971).
Each proper scoring rule induces a Bregman divergence, measuring the loss from incorrect prediction. The Bregman divergence is defined with the convex utility function of a scoring rule. See Figure 2 for a geometric demonstration of proper scoring rules.
Theorem 3.4 (McCarthy (1956); Savage (1971)).
A scoring rule is proper if and only if there exists a convex utility function $u:\Delta(\Theta)\to\mathbb{R}$ such that
$S(p,\theta)=u(p)+\nabla u(p)\cdot(\theta-p).$ (11)
Notice that when the prediction is the true distribution, ${\mathbf{E}}_{\theta\sim p}\left[S(p,\theta)\right]=u(p)$ since ${\mathbf{E}}\left[\theta\right]=p$.
Definition 3.5 (Bregman divergence).
Specified by a convex function $u:\Delta(\Theta)\to\mathbb{R}$, the Bregman divergence $\textsc{Breg}(p^{\prime},p)$ is defined as [^8] The definition of Bregman divergence here reverses the order of input in the conventional definition to align with proper scoring rules. In the conventional definition, $\textsc{Breg}(p,\widehat{p})=u(p)-u(\widehat{p})+\nabla u(\widehat{p})(\widehat{p}-p)$.
$\textsc{Breg}(p,\widehat{p})=u(\widehat{p})-u(p)+\nabla u(p)(p-\widehat{p}).$ (12)
Given a proper scoring rule $S$, we write the Bregman divergence defined with $u(p)={\mathbf{E}}_{\theta\sim p}\left[S(p,\theta)\right]$ as $\textsc{Breg}_{S}$. Breaking down CFDL to each prediction value in $Q$, it follows that the contribution of $q_{i}\in Q$ is:
$\frac{1}{n_{i}}\sum_{t\in[T]}(S(\widehat{q}_{i},\theta_{t})-S(q_{i},\theta_{t}))\mathbb{I}\left[p_{t}=q_{i}\right]={\mathbf{E}}_{\theta\sim\widehat{q}_{i}}\left[S(\widehat{q}_{i},\theta)-S(q_{i},\theta)\right]=\textsc{Breg}_{S}(q_{i},\widehat{q}_{i}).$
$p$$S(p,0)$$S(p,1)$$u(p)$$\widehat{p}$$u(\widehat{p})$$\textsc{Breg}(p,\widehat{p})$$S$001111111100state; reportscore
Figure 2: The graphic explanation of the connection between proper scoring rule and Bregman divergence. The thick convex curve plots the convex utility function $u(p)$ for a proper scoring rule. Fix a report, the score $S(p,\theta)=u(p)+\nabla u(p)(\theta-p)$ is the extreme points on the gradient hyperplane passing $u(p)$ (the thin line). Given empirical distribution $\widehat{p}$, the Bregman divergence $\textsc{Breg}(p,\widehat{p})$ is the loss of reporting $p$ instead of $\widehat{p}$.
Proposition 3.6.
Given a sequence of $T$ predictions $\bm{p}=(p_{t})_{t\in[T]}$ from a finite set $Q=\{q_{1},\ldots,q_{m}\}\subseteq[0,1]$ and realizations $\bm{\theta}=(\theta_{t})_{t\in[T]}$ of states, define $n_{i}$ and $\widehat{q}_{i}$ as in (1) and (2). For any proper scoring rule $S$,
$\textsc{CFDL}_{S}(\bm{p},\bm{\theta})=\frac{1}{T}\sum_{i\in[m]}n_{i}\textsc{Breg}_{S}(q_{i},\widehat{q}_{i}).$ (13)
3.2 Approximation via V-Bregman Divergences
Following ideas from Li ⓡ al. (2022); Kleinberg et al. (2023), we decompose any Bregman divergence into a linear combination on a basis. Thus, our CDL is bounded by the worst case CFDL on the basis, the V-Bregman divergences.
Definition 3.7 (V-Bregman divergence).
The V-Bregman divergence with kink $\mu$ is defined as
$\textsc{VBreg}_{\mu}(q,\widehat{q})=\begin{cases}0,&\text{if 1) }q<\mu\text{ and }\widehat{q}\leq\mu\text{ or 2)}q\geq\mu\text{ and }\widehat{q}\geq\mu;\\ \frac{|\widehat{q}-\mu|}{\max\{1-\mu,\mu\}},&\text{otherwise}.\end{cases}$ (14)
We note that any V-Bregman divergence can be induced by a V-shaped proper scoring rule bounded in $[0,1]$. More specifically, the V-shaped scoring rule offers two actions for the agent to decide, with optimal decision rule as a threshold at $\mu$. Without loss of generality, suppose $\mu\leq\nicefrac{{1}}{{2}}$. $\textsc{VBreg}_{\mu}$ can be induced from the following proper scoring rule, which is shown in Figure 3.
$S_{\mu}(p,\theta)=\left\{\begin{array}[]{cc}\nicefrac{{1}}{{2}}-\frac{1}{2}\cdot\frac{\theta-\mu}{1-\mu}&\text{if }p\leq\mu\\ \nicefrac{{1}}{{2}}+\frac{1}{2}\cdot\frac{\theta-\mu}{1-\mu}&\text{else},\end{array}\right.$ (15)
[…]
Figure 3: The thick black line plots the special convex utility function $u$ for the scoring rule. The convex utility function is V-shaped, consisting of two linear pieces intersecting at $\mu$. Once fixing the prediction $p$, the score $S(p,\theta)=u(p)+\nabla u(p)\cdot(\theta-p)$ is linear in the state $\theta$. The scoring rule offers two set of scores for the prediction to selecttwo linear lines for prediction. When $p\leq q_{0}$, the prediction selects scores $\{S(0,0),S(0,1)\}$. Otherwise, the prediction selects $\{S(1,0),S(1,1)\}$.
The following lemma allows us to decompose a general Bregman divergence as a linear combination of V-Bregman divergences.
Lemma 3.8 (Li ⓡ al. (2022); Kleinberg et al. (2023)).
Let $u:[0,1]\to[0,1]$ be a twice differentiable convex function. Let $S$ be the proper scoring rule generated by $u$ as in (11). Then for every $p,\widehat{p}\in[0,1]$,
$\textsc{Breg}_{S}(p,\widehat{p})=\int_{\mu=0}^{1}u^{\prime\prime}(\mu)\cdot\max(1-\mu,\mu)\textsc{VBreg}_{\mu}\text{d}\mu.$ (16)
Definition 3.9 (VCFDL).
Given a sequence of $T$ predictions $\bm{p}=(p_{t})_{t\in[T]}$ from a finite set $Q=\{q_{1},\ldots,q_{m}\}\subseteq[0,1]$ and realizations $\bm{\theta}=(\theta_{t})_{t\in[T]}$ of states, the VCFDL is defined as the supremem over all V-Bregman divergences:
$\textsc{VCFDL}=\sup_{\mu\in[0,1]}\frac{1}{T}\sum_{i\in[m]}n_{i}\textsc{VBreg}_{\mu}(q_{i},\widehat{q}_{i}).$ (17)
Theorem 3.10.
VCFDL is a constant-factor approximation of CDL. That is, for any sequence of predictions $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T}$ and any sequence of states $\bm{\theta}=(\theta_{1},\ldots,\theta_{T})\in\{0,1\}^{T}$,
$\textsc{VCFDL}(\bm{p},\bm{\theta})\leq\textsc{CDL}(\bm{p},\bm{\theta})\leq 2\textsc{VCFDL}(\bm{p},\bm{\theta}).$
Proof.
Since $\textsc{VBreg}_{\mu}$ is a special case of a Bregman divergence $\textsc{Breg}_{S}$ for a scoring rule $S$, comparing the definition of VCFDL in (17) and the expression of PSwap in (13), we immediately get $\textsc{VCFDL}(\bm{p},\bm{\theta})\leq\textsc{CDL}(\bm{p},\bm{\theta})$.
Now let us consider any proper scoring rule $S:[0,1]\times\{0,1\}\to[0,1]$. By Theorem 3.4, there exists a convex function $u:[0,1]\to[0,1]$ with bounded sub-gradients $\nabla u(p)\in[-1,1]$ such that (11) holds. We can construct a sequence of twice-differentiable convex functions $u_{1},u_{2},\ldots$ such that for every $p\in[0,1]$, $u_{j}(p)\to u(p)$ and $u_{j}^{\prime}(p)\to\nabla u(p)$ as $j\to+\infty$. Let $S_{j}$ denote the proper scoring rule generated by $u_{j}$ as in (11). For every $p,\widehat{p}\in[0,1]$, we have
$\lim_{j\to\infty}\textsc{Breg}_{S_{j}}(p,\widehat{p})=\textsc{Breg}_{S}(p,\widehat{p}).$
Let $Q=\{q_{1},\ldots,q_{m}\}\subseteq[0,1]$ be a finite prediction space such that $p_{t}\in Q$ for every $t\in[T]$. Define $n_{i}$ and $\widehat{q}_{i}$ as in (1) and (2). By Lemma 3.8,
$\sum_{i\in[m]}n_{i}\textsc{Breg}_{S_{j}}(q_{i},\widehat{q}_{i})$ $=\sum_{i\in[m]}n_{i}\int_{\mu=0}^{1}u^{\prime\prime}_{j}(\mu)\cdot\max(1-\mu,\mu)\textsc{VBreg}_{\mu}(q_{i},\widehat{q}_{i})\text{d}\mu$
$=\int_{\mu=0}^{1}u^{\prime\prime}_{j}(\mu)\cdot\max(1-\mu,\mu)\sum_{i\in[m]}n_{i}\textsc{VBreg}_{\mu}(q_{i},\widehat{q}_{i})\text{d}\mu$
$\leq\int_{\mu=0}^{1}u^{\prime\prime}_{j}(\mu)\cdot\max(1-\mu,\mu)\textsc{VCFDL}(\bm{p},\bm{\theta})\text{d}\mu$
$=\textsc{VCFDL}(\bm{p},\bm{\theta})\int_{\mu=0}^{1}u^{\prime\prime}_{j}(\mu)\cdot\max(1-\mu,\mu)\text{d}\mu$
$\leq\textsc{VCFDL}(\bm{p},\bm{\theta})\int_{\mu=0}^{1}u^{\prime\prime}_{j}(\mu)\cdot\text{d}\mu$
$=\textsc{VCFDL}(\bm{p},\bm{\theta})(u_{j}^{\prime}(1)-u_{j}^{\prime}(0)).$
The right hand side converges as $j\to\infty$:
$\lim_{j\to\infty}\textsc{VCFDL}(\bm{p},\bm{\theta})(u_{j}^{\prime}(1)-u_{j}^{\prime}(0))=\textsc{VCFDL}(\bm{p},\bm{\theta})(\nabla u(1)-\nabla u(0))\leq 2\textsc{VCFDL}(\bm{p},\bm{\theta}).$
Therefore, by Proposition 3.6,
$\textsc{CFDL}_{S}(\bm{p},\bm{\theta})=\sum_{q_{i}\in Q}n_{i}\textsc{Breg}_{S}(q_{i},\widehat{q}_{i})$ $=\lim_{j\to\infty}\sum_{q_{i}\in Q}n_{i}\textsc{Breg}_{S_{j}}(q_{i},\widehat{q}_{i})$
$\leq\lim_{j\to\infty}\textsc{VCFDL}(\bm{p},\bm{\theta})(u_{j}^{\prime}(1)-u_{j}^{\prime}(0))$
$\leq 2\textsc{VCFDL}(\bm{p},\bm{\theta}).$
Taking supremum over $S$ proves $\textsc{CDL}(\bm{p},\bm{\theta})\leq 2\textsc{VCFDL}(\bm{p},\bm{\theta})$. ∎
3.3 Computation of CDL
We allow the state space to be non-binary in this section. By solving a linear program, the CDL can be computed in time polynomial in the size of the prediction space $|Q|$ and the state space $|\Theta|$.
Theorem 3.11.
Given a sequence of predictions $\bm{p}$ and states $\bm{\theta}$, suppose $Q$ is the space of predictions. CDL can be computed in time polynomial in $|Q|$ and $|\Theta|$.
Proof.
The computation of CDL is an optimization problem over the space of proper scoring rules. We follow the idea in Li ⓡ al. (2022); Kleinberg et al. (2023). The optimal scoring rule can be computed by solving a linear program. Let $\widehat{Q}=\{\widehat{q}_{i}\}_{i}$ be the set of empirical distributions for each $q_{i}\in Q$. Define $\widehat{\Theta}=\{\tilde{\theta}\mid j\in|\Theta|,\tilde{\theta}_{j}=1,\tilde{\theta}_{j^{\prime}\neq j}=0\}$ as the set of indicator predictions of a certain state. Define the space of predictions as $\mathcal{Q}=Q\cup\widehat{Q}\cup\widehat{\Theta}$. We set the scores $s_{q,\theta},\forall q\in\mathcal{Q},\theta\in\Theta$ as variables in the linear program.
$\max_{S}\qquad$ $\sum_{q\in Q,\theta\in\Theta}(s_{\widehat{q},\theta}-s_{q,\theta})\widehat{q}(\theta)$
s.t. $s_{q,\theta}\in[0,1],\quad\forall q\in\mathcal{Q},\theta\in\Theta$ (bounded payoff)
$\sum_{\theta}q(\theta)s_{q,\theta}\geq\sum_{\theta}q(\theta)s_{q^{\prime},\theta},\quad\forall q,q^{\prime}\in\mathcal{Q}$ (properness)
The following proper scoring rule achieves the worst case CFDL.
$S(p,\theta)=s_{q,\theta},\text{ where }q=\operatorname*{arg\,max}_{q^{\prime}\in\mathcal{Q}}{\mathbf{E}}_{\theta\sim p}\left[s_{q^{\prime},\theta}\right].\qed$
4 CDL and Calibration Errors
In this section, we discuss connections of our CDL to calibration errors in the literature. We show that both ECE and $K_{2}$ are polynomially related to CDL (Theorem 4.1), but neither is a constant-factor approximation (Example 4.2). We give examples where the smooth calibration error differs significantly from CDL in either direction (Example 4.3). We also show the U-calibration error lowerbounds CDL (Proposition 4.4), but is not polynomially related (Example 4.5).
Theorem 4.1.
Given samples of predictions and corresponding states,
$\textsc{ECE}^{2}\leq$ $\textsc{CDL}\leq 2\textsc{ECE},$
$K_{2}\leq$ $\textsc{CDL}\leq 2\sqrt{K_{2}}.$
Proof of Theorem 4.1.
On the upper bound side, by 2.18, any proper scoring rule $S$ with range bounded in $[0,1]$ satisfies $\textsc{CFDL}_{S}\leq 2\textsc{ECE}$, implying $\textsc{CDL}\leq 2\textsc{ECE}$. It is easy to check that $\textsc{ECE}\leq\sqrt{K_{2}}$ (see e.g. Kleinberg et al., 2023; Roth, 2022). Thus we get the other upper bound $\textsc{CDL}\leq 2\sqrt{K_{2}}$.
On the lower bound side, by Lemma 2.19, we have $\textsc{CDL}\geq K_{2}$. Combining this with the fact $\textsc{ECE}\leq\sqrt{K_{2}}$ yields $\textsc{CDL}\geq\textsc{ECE}^{2}$.∎
All four inequalities in Theorem 4.1 are tight up to constant factors. We demonstrate the tight examples in Example 4.2.
Example 4.2.
Consider the following two miscalibrated predictors.
  • (a)
    (Tight example for upper bounds of CDL) The state is deterministically $1$. The predictor deterministically predicts $1-\epsilon$.
    In this case, $\textsc{ECE}=\epsilon$, $K_{2}=\epsilon^{2}$, $\textsc{CDL}=\Theta(\epsilon)$.
  • (b)
    (Tight example for lower bounds of CDL) The $T$ rounds are divided into $\sqrt{T}$ periods, each with $\sqrt{T}$ rounds. In each period $i$, the empirical distribution of the state is $\frac{i}{\sqrt{T}}$, and the predictor predicts $\frac{i}{\sqrt{T}}+\frac{1}{\sqrt{T}}$.
    In this case, $\textsc{ECE}=\frac{1}{\sqrt{T}}$, $K_{2}=\frac{1}{T}$, $\textsc{CDL}\in[\frac{1}{T},\frac{8}{T}]$.
From Example 4.2, we see that when predictions concentrate in a small interval, ECE calculates the CDL in the correct order as in (a). However, if predictions have high variance as in (b), the calibration error does not simply add up to the total loss in decision. Consider the V-Bregman divergence which corresponds to a decision problem with two actions. Suppose the kink, also the decision threshold, is at $\nicefrac{{1}}{{2}}$. Miscalibration at extreme predictions near 00 or $1$ will not induce a CFDL to the agent. This intuition is explained by Example 4.2, with which we prove the $\tilde{O}(\frac{1}{\sqrt{T}})$ CFDL later. We state it here and prove it in Section 5.1.
Define the bias in bucket $i$ (i.e. conditional ECE calibration error on prediction $q_{i}$):
$G_{i}=n_{i}|q_{i}-\widehat{q}_{i}|.$
{restatable*}
lemmalemSwapG Let $T,m$ be positive integers. Define $Q=\{q_{1},\ldots,q_{m}\}\subseteq[0,1]$ where $q_{i}=i/m$ for every $i=1,\ldots,m$. Given a sequence of predictions $\bm{p}=(p_{1},\ldots,p_{T})\in Q^{T}$ and realized states $\bm{\theta}=(\theta_{1},\ldots,\theta_{T})\in\{0,1\}^{T}$, define $n_{i}$ and $\widehat{q}_{i}$ as in (1) and (2). Define $G_{i}:=n_{i}|\widehat{q}_{i}-q_{i}|$. Fix a V-Bregman divergence with kink $\mu$, the CFDL is bounded by
$\textsc{CFDL}_{\mu}(\bm{p},\bm{\theta})\leq\frac{2}{T}\sum_{i\in[m]}\left(G_{i}-n_{i}|q_{i}-\mu|\right)_{+}.$
Proof of Example 4.2.
The calculation of ECE and $K_{2}$ are straightforward. We show the calculation of MSR separately for (a) and (b).
  • (a)
    By Example 4.2, $\textsc{CDL}\leq 2\textsc{VCFDL}\leq 4\epsilon$. We can find a Bregman divergence such that $\textsc{CFDL}\geq\frac{\epsilon}{2}$. Consider the scoring rule $S$ that has a V-Bregman divergence with kink $\mu=1-\frac{\epsilon}{2}$.
    $\textsc{CFDL}=\frac{|1-\mu|}{\mu}\geq\frac{\epsilon}{2}.$
    Thus, $\textsc{CDL}=\Theta(\epsilon)$.
  • (b)
    We can prove $\textsc{VCFDL}=\Theta(1)$. Fix any V-Bregman divergence with kink $\mu$, consider the corresponding $\textsc{CFDL}_{\mu}$.
    $\textsc{CFDL}_{\mu}$ $=\frac{1}{T}\cdot\sqrt{T}\sum_{i=1}^{\sqrt{T}}\frac{|\frac{i}{\sqrt{T}}-\mu|}{\max\{\mu,1-\mu\}}\left(\mathbb{I}\left[\frac{i}{\sqrt{T}}>\mu>\frac{i+1}{\sqrt{T}}\right]+\mathbb{I}\left[\frac{i}{\sqrt{T}}<\mu<\frac{i+1}{\sqrt{T}}\right]\right)$
    We notice that for predictions that induces a non-zero CFDL, it must be $|\frac{i}{\sqrt{T}}-\mu|\leq\frac{1}{\sqrt{T}}$. Since $\max\{\mu,1-\mu\}\geq\frac{1}{2}$,
    $\textsc{CFDL}_{\mu}\leq\frac{1}{\sqrt{T}}\cdot 2\cdot\frac{1}{\nicefrac{{1}}{{2}}}\frac{1}{\sqrt{T}}=\frac{4}{T}.$
    We know $\textsc{CDL}\leq 2\textsc{VCFDL}\leq\frac{8}{T}$.
CDL is not polynomially related to smCal or distCal. Specifically, example (b) in Example 4.3 shows the $\tilde{O}(\sqrt{T})$ distCal guarantee in online calibration does not apply to CDL.
Example 4.3.
We give two examples of miscalibrated predictors.
  • (a)
    (Large distCal, small CDL) The same example as (b) in Example 4.2.
    $\textsc{distCal}\geq\textsc{smCal}\geq\frac{1}{\sqrt{T}}$, $\textsc{CDL}\in[\frac{1}{T},\frac{2}{T}]$.
  • (b)
    (Small distCal, large CDL) At the first $\frac{T}{2}$ rounds, $\theta=1$ deterministically, and the predictor predicts $\nicefrac{{1}}{{2}}+\epsilon$. At the later $\frac{T}{2}$ rounds, $\theta=0$ deterministically, and the predictor predicts $\nicefrac{{1}}{{2}}-\epsilon$.
    $\textsc{smCal}\leq\textsc{distCal}\leq\epsilon$, $\textsc{CDL}=\Omega(1)$.
    We can take $\epsilon=\frac{1}{\sqrt{T}}$, which can be arbitrarily small.
Proof of Example 4.3.
We calculate distCal and CDL separately for each example.
  • (a)
    It only remains to show $\textsc{smCal}\geq\frac{1}{\sqrt{T}}$. By Definition 2.5, take Lipschitz function $\sigma(\cdot)=1$.
    $\textsc{smCal}\geq\frac{1}{T}\sum_{t}(p_{t}-\theta_{t})=\frac{1}{T}\cdot T\cdot\frac{1}{\sqrt{T}}=\frac{1}{\sqrt{T}}.$
  • (b)
    For distCal, a calibrated predictor always predicts $\nicefrac{{1}}{{2}}$.
    $\textsc{distCal}\leq\frac{1}{T}\sum_{t\in[T]}|p_{t}-\frac{1}{2}|=\epsilon.$
    For CDL, consider the V-Bregman divergence with kink $\nicefrac{{1}}{{2}}+2\epsilon$. The CFDL for this V-Bregman divergence is
    $\textsc{CFDL}_{\nicefrac{{1}}{{2}}}\geq\frac{1}{2}(1-\nicefrac{{1}}{{2}}-2\epsilon)=\nicefrac{{1}}{{4}}-\epsilon.$
    $\textsc{CDL}\geq\textsc{CFDL}_{\nicefrac{{1}}{{2}}}$ implies $\textsc{CDL}=\Omega(1)$.
Vanishing U-calibration error is necessary but not sufficient for calibration. By definition, the U-calibration error lowerbounds CDL.
Proposition 4.4.
For any sequence of predictions and states,
[…]
Kleinberg et al. (2023) gives an example where in the limit as $T\to\infty$, the U-calibration error is 00, while ECE, $K_{2}$ and smCal are non-zero. We present a simpler example showing $\textsc{UCal}=0$ is insufficient for calibration and $\textsc{CDL}=\Omega(1)$.
Example 4.5.
The empirical distribution of the state is $\Pr[\theta=1]=\nicefrac{{1}}{{2}}$, i.e. $\frac{T}{2}$ samples are $1$, $\frac{T}{2}$ samples are 00. The predictor predicts $\frac{3}{4}$ when the state is $1$, and $\frac{1}{4}$ when the state is 00.
In this example, $\textsc{UCal}=0$, $\textsc{CDL}=\Omega(1)$.
Proof of Example 4.5.
First, UCal is always non-negative. By definition, there exists a degenerate decision task with a constant payoff, where the decision maker has 00 external regret. By the same V-shaped decomposition in Lemma 3.8 and in Kleinberg et al. (2023), it only remains to show this predictor is weakly better than predicting $\frac{1}{2}$ on all V-shaped scoring rules in Equation 15.
Fix each V-shaped scoring rule with kink $\mu\in[0,\frac{1}{4}]$, always predicting $\nicefrac{{1}}{{2}}$ and the miscalibrated predictor aboth achieve the same payoff. The external regret is thus 00. For V-shaped scoring rules with kink $\mu\in(\frac{1}{4},\frac{1}{2}]$, predicting $\frac{1}{2}$ yields payoff $\frac{3}{4}-\frac{1}{4}\cdot\frac{\mu}{1-\mu}$. However, the miscalibrated predictor obtains higher payoff $\frac{3}{4}+\frac{1}{4}\cdot\frac{\mu}{1-\mu}$, which achieves negative external regret. The case for $\mu>\frac{1}{2}$ is similar to the two cases above. We can conclude that the external regret for each downstream decision task is non-positive.
This example, however, is very miscalibrated. Specifically, consider the V-shaped scoring rule with kink at $\frac{1}{4}-\epsilon$. CFDL of the miscalibrated predictor is $\Omega(1)$ for predicting $\frac{1}{4}$, where the conditional empirical frequency is 00.
5 Minimizing Calibration Decision Loss
In this section, we present our online binary prediction algorithm (Algorithm 1) and prove that it achieves the following low CDL guarantee:
Theorem 5.1.
For $T\geq 2$, Algorithm 1 runs in time polynomial in $T$ and makes predictions $\bm{p}=(p_{1},\ldots,p_{T})$ satisfying
${\mathbf{E}}\left[\textsc{CDL}(\bm{p},\bm{\theta})\right]\leq O(\frac{\log T}{\sqrt{T}}).$
Here, $\bm{\theta}=(\theta_{1},\ldots,\theta_{T})$ is the sequence of realized states chosen by any adversary in the online binary prediction setting (see Section 2.5), and the expectation is over the randomness of the algorithm.
We design Algorithm 1 such that it makes predictions in a finite set $Q:=\{q_{1},\ldots,q_{m}\}\subseteq[0,1]$, where $q_{i}=i/m$ for each $i=1,\ldots,m$. Later we will pick the optimal choice of $m\approx\sqrt{T}/\log T$. We view each $q_{i}$ as a bucket, so the prediction $p_{t}$ made by Algorithm 1 in each round $t$ falls into one of the $m$ buckets $q_{1},\ldots,q_{m}$. We use $n_{i}$ to denote the number of predictions in bucket $i$ (see (1)), and use $\widehat{q}_{i}$ to denote the average value of the realized states corresponding to the $n_{i}$ predictions (see (2)). We define $G_{i}:=n_{i}|q_{i}-\widehat{q}_{i}|$ as the bias from bucket $i$.
In Section 5.1, we prove a key technical lemma which allows us to attribute the CDL to the bias $G_{i}$ from each bucket. We then present Algorithm 1 in Section 5.2 and complete the proof of Theorem 5.1.
5.1 Attributing CDL to Bucket-wise Biases
We establish our key technical lemma that allows us to upper bound CDL using the bucket-wise biases.
Lemma 5.2 (Formal and generalized version of Lemma 1.3).
Let $T,m\geq 2$ be positive integers. Define $Q=\{q_{1},\ldots,q_{m}\}\subseteq[0,1]$ where $q_{i}=i/m$ for every $i=1,\ldots,m$. Given a sequence of predictions $\bm{p}=(p_{1},\ldots,p_{T})\in Q^{T}$ and realized states $\bm{\theta}=(\theta_{1},\ldots,\theta_{T})\in\{0,1\}^{T}$, define $n_{i}$ and $\widehat{q}_{i}$ as in (1) and (2). Define $G_{i}:=n_{i}|\widehat{q}_{i}-q_{i}|$. For $\alpha,\beta\geq 0$, define maximum deviation $\mathcal{D}$:
$\mathcal{D}(\bm{p},\bm{\theta}):=\max_{1\leq i\leq m}\{[G_{i}-\alpha\sqrt{n}_{i}-\beta n_{i}]_{+}\},\quad\text{where $[x]_{+}:=\max(x,0)$.}$ (18)
Then
$\textsc{CDL}(\bm{p},\bm{\theta})\leq 4\frac{m}{T}\mathcal{D}(\bm{p},\bm{\theta})+\frac{4\alpha}{\sqrt{T}}+4\beta+O(\frac{\alpha^{2}m\log m}{T}).$
Our proof of Lemma 5.2 relies on the following helper lemma which controls the CFDL w.r.t. a single V-shaped Bregman divergence.
[…]
Proof.
$\textsc{CFDL}_{\mu}(\bm{p},\bm{\theta})$ $=\frac{1}{T}\sum_{i\in[m]}n_{i}\textsc{VBreg}_{\mu}(q_{i},\widehat{q}_{i})$
$=\frac{1}{T}\sum_{i\in[m]}n_{i}\frac{|\widehat{q}_{i}-\mu|}{\max\{1-\mu,\mu\}}\left(\mathbb{I}\left[q_{i}<\mu<\widehat{q}_{i}\right]+\mathbb{I}\left[q_{i}>\mu>\widehat{q}_{i}\right]\right)$ (by (14))
$\leq\frac{1}{T}\sum_{i\in[m]}n_{i}\frac{\left(|\widehat{q}_{i}-q_{i}|-|\mu-q_{i}|\right)_{+}}{\max\{1-\mu,\mu\}}$
$\leq\frac{2}{T}\sum_{i\in[m]}\left(G_{i}-n_{i}|q_{i}-\mu|\right)_{+}.\qed$
Proof of Lemma 5.2.
By Theorem 3.10, it suffices to prove
$\textsc{VCFDL}(\bm{p},\bm{\theta})\leq 2\frac{m}{T}\mathcal{D}(\bm{p},\bm{\theta})+\frac{2\alpha}{\sqrt{T}}+2\beta+O(\frac{\alpha^{2}m\log m}{T}).$ (19)
For any $\mu\in[0,1]$, by Example 4.2,
$\textsc{CFDL}_{\mu}(\bm{p},\bm{\theta})$ $\leq\frac{2}{T}\sum_{i=1}^{m}(G_{i}-n_{i}|q_{i}-\mu|)_{+}$
$\leq\frac{2}{T}\sum_{i=1}^{m}(\mathcal{D}(\bm{p},\bm{\theta})+\alpha\sqrt{n}_{i}+\beta n_{i}-n_{i}|q_{i}-\mu|)_{+}$
$\leq\frac{2}{T}\sum_{i=1}^{m}(\mathcal{D}(\bm{p},\bm{\theta})+\beta n_{i}+(\alpha\sqrt{n_{i}}-n_{i}|q_{i}-\mu|)_{+})$ (because $\mathcal{D}(\bm{p},\bm{\theta})\geq 0$ and $\beta\geq 0$)
$=\frac{2m}{T}\mathcal{D}(\bm{p},\bm{\theta})+2\beta+\frac{2}{T}\sum_{i=1}^{m}(\alpha\sqrt{n_{i}}-n_{i}|q_{i}-\mu|)_{+}.$ (20)
Let us re-arrange $q_{1},\ldots,q_{m}$ in non-decreasing order of $|q_{i}-\mu|$. That is, we choose a bijection $\tau$ from $\{1,\ldots,m\}$ to itself such that $|q_{\tau(i)}-\mu|$ is a non-decreasing function of $i$. When $i=1$, we use the following trivial upper bound:
$(\alpha\sqrt{n_{\tau(1)}}-n_{\tau(1)}|q_{\tau(1)}-\mu|)_{+}\leq\alpha\sqrt{n_{\tau(1)}}\leq\alpha\sqrt{T}.$ (21)
When $i>1$, we have $|q_{\tau(i)}-\mu|\geq\Omega(i/m)$, and thus
$(\alpha\sqrt{n_{\tau(i)}}-n_{\tau(i)}|q_{\tau(i)}-\mu|)_{+}\leq\frac{\alpha^{2}}{4|q_{\tau(i)}-\mu|}=O(\alpha^{2}m/i).$ (22)
Plugging (21) and (22) into (20), we get
$\textsc{CFDL}_{\mu}(\bm{p},\bm{\theta})$ $\leq\frac{2m}{T}\mathcal{D}(\bm{p},\bm{\theta})+2\beta+\frac{2\alpha}{\sqrt{T}}+O\left(\frac{1}{T}\alpha^{2}m\sum_{i=2}^{m}\frac{1}{i}\right)$
$\leq 2\frac{m}{T}\mathcal{D}(\bm{p},\bm{\theta})+\frac{2\alpha}{\sqrt{T}}+2\beta+O(\frac{\alpha^{2}m\log m}{T}).$
This implies (19), as desired. ∎
5.2 Efficient CDL Minimization Algorithm
Given Lemma 5.2, we can establish the low CDL guarantee in Theorem 5.1 by designing an algorithm (Algorithm 1) that minimizes $\mathcal{D}(\bm{p},\bm{\theta})$. Specifically, for parameters $m\approx\sqrt{T}/\log T,\beta=1/m,\alpha\approx\sqrt{\log T}$, we show that Algorithm 1 achieves ${\mathbf{E}}\left[\mathcal{D}(\bm{p},\bm{\theta})\right]=O(\log T)$ (Lemma 5.5). Our design of Algorithm 1 largely follows the ideas from Noarov et al. (2023), but we make small but important refinements to obtain a stronger guarantee as needed to prove Theorem 5.1 (see Remark 5.6).
In Algorithm 1, we partition the interval $[0,1]$ into $m$ sub-intervals $I_{1},\ldots,I_{m}$ where
$I_{1}=[0,1/m],I_{2}=(1/m,2/m],\ldots,I_{m}=((m-1)/m,1].$ (23)
In each round $t=1,\ldots,T$, Algorithm 1 first computes a prediction $\widetilde{p}_{t}\in[0,1]$ and then outputs a discretized prediction $p_{t}$ via rounding. Specifically, $Q=\{q_{1},\ldots,q_{m}\}$ is the discretized prediction space, where $q_{i}=i/m$ for every $i=1,\ldots,m$. The prediction $\widetilde{p}_{t}$ belongs to an interval $I_{i}$ for a unique index $i=1,\ldots,m$, and the corresponding discretized prediction is $p_{t}=q_{i}\in Q$. We use $n_{i}$ to denote the number of rounds $t$ in which $p_{t}=q_{i}$, or equivalently, $\widetilde{p}_{t}\in I_{i}$:
$n_{i}:=\sum_{t=1}^{T}\mathbb{I}\left[p_{t}=q_{i}\right]=\sum_{t=1}^{T}\mathbb{I}\left[\widetilde{p}_{t}\in I_{i}\right].$
For each $i=1,\ldots,m$, and $\sigma=\pm 1$, we define
$l_{i,\sigma}(p,\theta):=\sigma\mathbb{I}\left[p\in I_{i}\right](p-\theta)\quad\text{for every prediction $p\in[0,1]$ and state $\theta\in\{0,1\}$.}$ (24)
Algorithm 1 calls an expert regret minimization oracle $\mathcal{A}$ from Chen et al. (2021). Here we imagine $12m+$ 1 experts: one expert for each pair $(i,\sigma)\in[m]\times\{\pm 1\}$ and one extra auxiliary expert. In each round $t$, the oracle $\mathcal{A}$ computes a distribution over the experts represented by values $w_{t,i,\sigma}\geq 0$, where for each pair $(i,\sigma)$, the value $w_{t,i,\sigma}\geq 0$ is the probability mass on the expert corresponding to $(i,\sigma)$. Thus, the probability mass on the auxiliary expert is $1-\sum_{(i,\sigma)}w_{t,i,\sigma}\geq 0$.
The distribution computed by $\mathcal{A}$ in each round $t$ is based on the gains $l_{t^{\prime},i,\sigma}\in[-1,1]$ received by each expert $(i,\sigma)$ at Step 7 in each previous round $t^{\prime}<t$. We set the gain of the auxiliary expert to always be zero. The work of Chen et al. (2021) shows a construction of the oracle with the following property:
Lemma 5.3 (Chen et al. (2021), applied to Algorithm 1).
For some absolute constant $C>0$, there exists an expert regret minimization oracle $\mathcal{A}$ for step 1 of Algorithm 1 with the following properties. Assume $T,m\geq 2$. In each round $t=1,\ldots,T$, the oracle computes $w_{t,i,\sigma}\geq 0$ for every $i\in[m]$ and $\sigma=\pm 1$ in time $\mathsf{poly}(m)$ such that
$\sum_{i\in[m],\sigma=\pm 1}w_{t,i,\sigma}$ $\leq 1,$
$-\sum_{t=1}^{T}\sum_{i\in[m],\sigma=\pm 1}{w_{t,i,\sigma}}l_{t,i,\sigma}$ $\leq C\log(mT),$ (25)
and for every $i=1,\ldots,m$ and $\sigma=\pm 1$,
$\sum_{t=1}^{T}l_{t,i,\sigma}-\sum_{t=1}^{T}\sum_{i^{\prime}\in[m],\sigma^{\prime}=\pm 1}{w_{t,i^{\prime},\sigma^{\prime}}}l_{t,i^{\prime},\sigma^{\prime}}\leq C\left(\log(mT)+\sqrt{n_{i}\log(mT)}\right).$ (26)
For each expert $(i,\sigma)$, the guarantee (26) is stronger than more standard guarantees for the experts problem in that the right hand side of (26) has an $\sqrt{n_{i}}$ dependence rather than an $\sqrt{T}$ dependence. For the auxiliary expert, we get the guarantee (25), which can be viewed as a special form of (26) with $l_{t,i,\sigma}=0$ and $n_{i}=0$.
Algorithm 1 Algorithm for MSR minimization.
Parameters: positive integers $m,T$; $\epsilon>0$; discretized prediction space $Q=\{q_{1},\ldots,q_{m}\}$ where $q_{i}=i/m$; intervals $I_{1},\ldots,I_{m}$ partitioning $[0,1]$ as defined in (23); functions $l_{i,\sigma}$ as defined in (24).
for each round $t=1,\ldots,T$ do
     1. Compute expert weights $w_{t,i,\sigma}$ for every $i\in[m]$ and $\sigma=\pm 1$ using the expert regret minimization oracle $\mathcal{A}$ from Lemma 5.3.
     2. For any distribution $s$ over $[0,1]$, define
$h_{t}(s):=\max_{\theta\in\{0,1\}}{\mathbf{E}}_{p\sim s}\left[\sum_{i\in[m],\sigma=\pm 1}w_{t,i,\sigma}l_{i,\sigma}(p,\theta)\right].$ (27)
     3. Find distribution $s_{t}$ such that $h_{t}(s_{t})\leq\epsilon$.
     4. Draw $\tilde{p}_{t}\in[0,1]$ from distribution $s_{t}$.
     5. Output $p_{t}:=q_{i}$, where $i$ is the unique index in $\{1,\ldots,m\}$ satisfying $\widetilde{p}_{t}\in I_{i}$.
     6. Receive the realized state $\theta_{t}\in\{0,1\}$.
     7. Calculate expert gains $l_{t,i,\sigma}$ for every $i\in[m]$ and $\sigma=\pm 1$:
$l_{t,i,\sigma}:=l_{i,\sigma}(\widetilde{p}_{t},\theta_{t}).$ (28)
end for
The following lemma shows that Step 3 of Algorithm 1 can be computed efficiently:
Lemma 5.4 (Noarov et al. (2023)).
Let $\epsilon>0$ be a parameter. At step 3 of Algorithm 1, a solution $s_{t}$ satisfying $h_{t}(s_{t})\leq\epsilon$ always exists and can be computed in time $\mathsf{poly}(\epsilon^{-1})$.
The existence of $s_{t}$ in Lemma 5.4 can be proved using the minimax theorem. We refer the reader to Noarov et al. (2023) for a complete proof of (a more general version of) Lemma 5.4 which includes an efficient algorithm for computing $s_{t}$ using the Follow-the-Perturbed-Leader approach.
We are now ready to prove that Algorithm 1 guarantees a small value of $\mathcal{D}(\bm{p},\bm{\theta})$ in expectation:
Lemma 5.5.
Let $C>0$ be the absolute constant from Lemma 5.3. Assume $m,T\geq 2$ and $\varepsilon=1/T$ in Algorithm 1. Let $\bm{p}=(p_{1},\ldots,p_{T})$ be the predictions made by Algorithm 1 on the adversarially chosen states $\bm{\theta}=(\theta_{1},\ldots,\theta_{T})$. Define $\mathcal{D}$ as in (18) for $\alpha=C\sqrt{\log(mT)},\beta=1/m$. Then,
${\mathbf{E}}\left[\mathcal{D}(\bm{p},\bm{\theta})\right]\leq O(\log(mT)),$
where the expectation is over the randomness of Algorithm 1.
Remark 5.6.
We define $\mathcal{D}(\bm{p},\bm{\theta})$ in (18) as a maximum over the buckets $i$. Thus, Lemma 5.5 shows an upper bound on the “expectation of maximum”, which is stronger than guarantee stated in Noarov et al. (2023) on the “maximum of expectation”. In our proof below, we use a slightly more careful analysis than what is used by Noarov et al. (2023) to achieve the stronger guarantee.
Proof of Lemma 5.5.
For every $i=1,\ldots,m$, by the definition of $G_{i}$ in Lemma 5.2, we have
$G_{i}=n_{i}|q_{i}-\widehat{q}_{i}|$ $=\left|\sum_{t=1}^{T}\mathbb{I}\left[p_{t}=q_{i}\right](p_{t}-\theta_{t})\right|$
$\leq\left|\sum_{t=1}^{T}\mathbb{I}\left[p_{t}=q_{i}\right](\widetilde{p}_{t}-\theta_{t})\right|+n_{i}/m$
$=\max_{\sigma=\pm 1}\sum_{t=1}^{T}l_{i,\sigma}(\widetilde{p}_{t},\theta_{t})+n_{i}/m$ (by (24))
$=\max_{\sigma=\pm 1}\sum_{t=1}^{T}l_{t,i,\sigma}+n_{i}/m.$ (by (28))
Therefore, by (26),
$G_{i}-n_{i}/m-\alpha\sqrt{n_{i}}\leq\left(\max_{\sigma=\pm 1}\sum_{t=1}^{T}l_{t,i,\sigma}\right)-\alpha\sqrt{n_{i}}\leq\sum_{t=1}^{T}\sum_{i^{\prime}\in[m],\sigma^{\prime}=\pm 1}{w_{t,i^{\prime},\sigma^{\prime}}}l_{t,i^{\prime},\sigma^{\prime}}+C\log(mT).$
By (25),
$0\leq\sum_{t=1}^{T}\sum_{i\in[m],\sigma=\pm 1}{w_{t,i,\sigma}}l_{t,i,\sigma}+C\log(mT).$
Combining the two inequalities above, we get
$\mathcal{D}(\bm{p},\bm{\theta})\leq\sum_{t=1}^{T}\sum_{i\in[m],\sigma=\pm 1}{w_{t,i,\sigma}}l_{t,i,\sigma}+C\log(mT).$ (29)
By the definition of $h_{t}$ in (27), the guarantee of $h_{t}(s_{t})\leq\epsilon$, and the fact that $\widetilde{p}_{t}$ is drawn from $s_{t}$, we get
${\mathbf{E}}\left[\sum_{t=1}^{T}\sum_{i\in[m],\sigma=\pm 1}{w_{t,i,\sigma}}l_{t,i,\sigma}\right]\leq\sum_{t=1}^{T}\epsilon\leq 1.$ (30)
Combining (29) and (30), we get
${\mathbf{E}}\left[\mathcal{D}(\tilde{\bm{p}},\bm{\theta})\right]\leq O(\log(mT)).\qed$
We now complete the proof of our main theorem.
Proof of Theorem 5.1.
We choose $\epsilon=1/T$ in Algorithm 1 and set $m=\Theta(\sqrt{T}/\log T)$. Following the setting of Lemma 5.5, we define $\mathcal{D}$ as in (18) for $\alpha=C\sqrt{\log(mT)},\beta=1/m$.
We have
${\mathbf{E}}\left[\textsc{CDL}(\bm{p},\bm{\theta})\right]$ $\leq\frac{4m}{T}{\mathbf{E}}\left[\mathcal{D}(\bm{p},\bm{\theta})\right]+\frac{4\alpha}{\sqrt{T}}+4\beta+O(\frac{\alpha^{2}m\log m}{T})$ (by Lemma 5.2)
$\leq O(\frac{m\log(mT)}{T})+\frac{4\alpha}{\sqrt{T}}+4\beta+O(\frac{\alpha^{2}m\log m}{T})$ (by Lemma 5.5)
$=O(\frac{m\log T}{T})+O(\frac{\sqrt{\log T}}{\sqrt{T}})+O(\frac{\log T}{\sqrt{T}})+O(\frac{\log T}{\sqrt{T}})$
$=O(\frac{\log T}{\sqrt{T}}).$
The running time guarantee of Algorithm 1 follows from Lemmas 5.3 and 5.4. ∎
Appendix A Minimax Proof for Minimizing CDL
We can prove the existence of an algorithm that achieves $\tilde{O}(\sqrt{T})$ CDL via the minimax theorem. The minimax theorem demonstrates a remarkable difference between $K_{1}$ and CDL - simply by reporting the adversary’s strategy truthfully, CDL matches the lowerbound.
In this setup, a predictor $F$ makes a prediction $p_{t}\in[0,1]$ at each time step $t=1,2,\ldots$, and an adversary $A$ picks an outcome $\theta_{t}\in\{0,1\}$. Both $p_{t}$ and $\theta_{t}$ are chosen based on the history $h_{t-1}=(p_{1},\theta_{1},p_{2},\theta_{2},\ldots,p_{t-1},\theta_{t-1})$, but these two quantities at the same time step $t$ cannot depend on each other. The goal of the predictor is to minimize $\textsc{CDL}(h_{T})$.
We can identify a predictor $F$ by its strategy at each time step. That is, we can write $F=(F_{1},\ldots,F_{T})$, where $F_{t}$ is a function mapping from $h_{t-1}$ to $p_{t}$. Similarly, we can identify the adversary $A$ by its strategy $A_{t}$ at each time step $t$, writing $A$ as $(A_{1},\ldots,A_{T})$. Given the strategies $F$ and $A$ of both players, we use $h_{F,A}$ to denote the history $h_{T}=(p_{1},\theta_{1},\ldots,p_{T},\theta_{T})$ generated by executing these strategies.
We will allow the predictor to randomize and play a mixed strategy which can be identified as a probability distribution ${\mathcal{F}}$ over all predictor strategies $F$.
Theorem A.1 (Existence).
There exists a mixed predictor strategy ${\mathcal{F}}$ such that
$\max_{A}\operatorname{\mathbb{E}}_{F\sim{\mathcal{F}}}[\textsc{CDL}(h_{F,A})]=O(\log(T)\sqrt{T}),$
where the maximum is over all strategies $A$ of the adversary.
In fact, we will prove a stronger version of Theorem A.1, where we restrict the predictor to make predictions $p_{t}$ in a discretized space $Q=\{q_{i}=\frac{i}{m}\}_{i\in[m]}\subseteq[0,1]$. We say a predictor strategy $F=(F_{1},\ldots,F_{T})$ is discretized if each $F_{t}$ is a function mapping from history $h_{t-1}$ to $p_{t}\in Q$. This discretization makes the (pure) strategy space of the predictor to be finite, allowing us to apply the minimax theorem.
We prove the following argument which implies Theorem A.1 immediately:
$\min_{{\mathcal{F}}}\max_{A}\operatorname{\mathbb{E}}_{F\sim{\mathcal{F}}}[\textsc{CDL}(h_{F,A})]=O(\log(T)\sqrt{T}),$
where the minimum is over a distribution ${\mathcal{F}}$ over discretized predictor strategies $F$, and the maximum is over all strategies $A$ of the adversary.
By the minimax theorem,
$\min_{{\mathcal{F}}}\max_{A}\operatorname{\mathbb{E}}_{F\sim{\mathcal{F}}}[\textsc{CDL}(h_{F,A})]=\max_{{\mathcal{A}}}\min_{F}\operatorname{\mathbb{E}}_{A\sim{\mathcal{A}}}[\textsc{CDL}(h_{F,A})].$
Thus, it suffices to reverse the order of play and prove Lemma A.2.
Lemma A.2.
The following holds:
$\max_{{\mathcal{A}}}\min_{F}\operatorname{\mathbb{E}}_{A\sim{\mathcal{A}}}[\textsc{CDL}(h_{F,A})]\leq O(\log(T)\sqrt{T}).$
where the minimum is over a deterministic discretized predictor strategy $F$, and the maximum is over a distribution over strategies $A$ of the adversary.
To show Lemma A.2, consider the truthful strategy of the predictor. At round $t$, let $\tilde{p}_{t}$ denote the conditional expectation of the adversary’s choice $\theta_{t}$ given past history. The predictor outputs the closest value $p_{t}$ in $Q$ to $\tilde{p}_{t}$.
Lemma A.3.
Let $\bm{p}=(p_{1},\ldots,p_{T})$ and $\bm{\theta}=(\theta_{1},\ldots,\theta_{T})$ be the sequence of predictions and states generated when the predictor uses the truthful strategy above. For $\alpha\geq 0$ and $\beta=\frac{1}{m}$, define $\mathcal{D}(\bm{p},\bm{\theta})$ as in Equation 18. We have
${\mathbf{E}}\left[\mathcal{D}(\bm{p},\bm{\theta})\right]\leq 2T^{2}m\exp{(-\alpha^{2}/2)}.$
Proof of Lemma A.3.
For bucket $i$, construct random variables $X_{1}^{(i)},\dots X_{T}^{(i)}$ such that $X_{j}^{(i)}=\theta_{t_{j}}-\widetilde{p}_{t_{j}}$, where $t_{j}$ is the index of the $j$-th round in which the predictor predicts $q_{i}$. If the number $n_{i}$ of rounds with prediction $q_{i}$ is smaller than $j$, we define $X_{j}^{(i)}=0$. For any $n\in[T]$, by Azuma’s inequality,
$\Pr\left[\left|\sum_{j=1}^{n}X_{j}^{(i)}\right|>\alpha\sqrt{n}\right]\leq 2\exp{(\frac{-\alpha^{2}}{2})}.$
By the union bound,
$\Pr\left[\exists n\in[T],i\in[m],\left|\sum_{j=1}^{n}X_{j}^{(i)}\right|>\alpha\sqrt{n}\right]\leq 2Tm\exp{(\frac{-\alpha^{2}}{2})}.$ (31)
Therefore, with probability at least $1-2Tm\exp{(\frac{-\alpha^{2}}{2})}$, for every $i\in[m]$, we have
$\left|\sum_{j=1}^{n_{i}}X_{j}^{(i)}\right|\leq\alpha\sqrt{n_{i}}.$
This implies
$G_{i}=\left|\sum_{j=1}^{n_{i}}(p_{t_{j}}-\theta_{t_{j}})\right|\leq\left|\sum_{j=1}^{n_{i}}(\widetilde{p}_{t_{j}}-\theta_{t_{j}})\right|+n_{i}/m\leq\alpha\sqrt{n_{i}}+\beta n_{i}.$
Thus, $\mathcal{D}(\bm{p},\bm{\theta})=0$ with probability at least $1-2Tm\exp{(\frac{-\alpha^{2}}{2})}$. Since $\mathcal{D}(\bm{p},\bm{\theta})\leq T$,
${\mathbf{E}}\left[\mathcal{D}(\bm{p},\bm{\theta})\right]\leq 2T^{2}m\exp{(\frac{-\alpha^{2}}{2})}.$
Proof of Lemma A.2.
Taking $\alpha=2\sqrt{\log(Tm)}$ in Lemma A.3, we get
${\mathbf{E}}\left[\mathcal{D}(\bm{p},\bm{\theta})\right]\leq 2.$
We choose $m=\Theta(\sqrt{T}/\log(T))$ buckets and set $\beta=\frac{1}{m}$. By Lemma 5.2,
${\mathbf{E}}\left[\textsc{CDL}(\bm{p},\bm{\theta})\right]$ $\leq 4m{\mathbf{E}}\left[\mathcal{D}(\bm{p},\bm{\theta})\right]+4\alpha\sqrt{T}+4\beta T+O(\alpha^{2}m\log m)$
$\leq O(\sqrt{T}/\log(T))+O(\sqrt{T\log(T)})+O(\sqrt{T}\log T)+O(\sqrt{T}\log(T))$
$=O(\sqrt{T}\log T).\qed$